### v3

# Fair Ordering via Streaming Social Choice Theory

Abstract.
Prior work (themis; kelkar2020order; cachin2022quick; vafadar2023condorcet) studies the question of “fairly” ordering transactions in a replicated state machine. Each of $n$ replicas receives transactions over a network in a possibly different order, and the system must aggregate these observed orderings into a single order. We argue that this problem is best viewed through the lens of social choice theory, in which (in the preference aggregation problem) rankings on candidates are aggregated into an election result.
Two features make this problem novel and distinct. First, the number of transactions is unbounded, and an ordering must be defined over a countably infinite set. And second, decisions must be made quickly and with only partial information. Additionally, some faulty replicas might alter their reported observations; their influence on the output should be bounded and well understood.
The prior work studies a “$\gamma$-batch-order-fairness” property, which divides an ordering into contiguous batches. If a $\gamma$ fraction of replicas receive a transaction $\tau$ before another transaction $\tau^{\prime}$, then $\tau^{\prime}$ cannot be in an earlier batch than $\tau$. We strengthen this definition to require that batches have minimal size ($\gamma$-batch-order-fairness can be vacuously satisfied by arbitrarily large batches) while accounting for the possibility of faulty replicas.
The lens of social choice theory allows us to give an ordering protocol with both strictly stronger fairness guarantees and better liveness properties than prior work. We study the Ranked Pairs method. Analysis of how missing information moves through the algorithm allows our streaming version to know when it can output a transaction. Deliberate construction of a tiebreaking rule ensures our algorithm outputs a transaction after a bounded time (in a synchronous network). Prior work relies on a fixed choice of $\gamma$ and bound on the number of faulty replicas $f$, but we show that Ranked Pairs satisfies our definition for every $\frac{1}{2}<\gamma\leq 1$ simultaneously and for any $f$, where fairness guarantees linearly degrade as $f$ increases.
copyright: none
1. Introduction
We study the problem of ordering transactions in the widely-used replicated state machine architecture (e.g., (ongaro2014search; lamport2001paxos; yin:hotstuff; oki1988viewstamped)). In the standard setting, each of a set of $n$ distinct replicas maintains a copy of a state machine. Replicas communicate to agree on a totally-ordered log of transactions $\tau_{1}\prec\tau_{2}\prec\tau_{3}\prec\dots$, and then each replica applies the transactions in this order to its local copy of the state machine. An (unbounded) set of clients create and broadcast new transactions. A classic example is a banking application, in which the state machine maintains a list of user account balances, and a transaction transfers money from one user to another.
This architecture requires that all replicas apply transactions in the same order. In the banking example, a client $A$ with a \$1 balance might create two transactions, one which sends a \$1 payment to client $B$ and one which sends a \$1 payment to client $C$. Only one of payments can execute successfully (assuming $A$’s balance is not allowed to overdraft) so if different replicas apply these transactions in different orders, then they will disagree on the balances of clients $B$ and $C$.
There are many communication protocols through which replicas can agree on some total order (solving the “total order broadcast” problem; broadcastsurvey gives a survey). Yet in many systems, most notably within today’s public blockchains, significant financial value can be derived from ordering transactions in specific ways (daian2020flash). These systems must therefore agree on not just any ordering but an “optimal” one, for some notion of optimal.
The key observation for this work is that this problem is a novel streaming variation of the classic preference aggregation problem of social choice theory (hagele2001lulls; llull; colomer2013ramon; condorcet1785essay; arrow1950difficulty). Prior work on this problem (kelkar2020order; themis; cachin2022quick) focuses on network communication, but (through varying protocols) each produces agreement on a “vote” from each replica on the ordering of a set of transactions (in social choice parlance, a “ranking” on a set of “candidates”). Agnostic to the choice of network protocol, how should a system aggregate a set of proposed orders (“rankings”) into a single, total order?
Two key differences separate this problem from the classic social choice setting. First, the number of transactions (“candidates”) is countably infinite (as clients can continually send new transactions for an unbounded time). And second, the system must produce the output ranking in a streaming fashion. It cannot wait to see the reported orders over the entire (infinite) set of transactions before making a decision on the relative ordering of two transactions. Instead, the system must produce as output an append-only, totally-ordered log of transactions (to be given to the state machine). Furthermore, the system should ideally minimize the delay between when a client sends a transaction and when that transaction is appended to the output log (i.e. the system needs “liveness”).
While we believe this problem to be interesting in its own right, studying it directly through the lens of classic social choice theory, enables us to develop an ordering algorithm with both much stronger order “fairness” guarantees (the key property studied in prior work, discussed below) and stronger liveness guarantees than all of the prior work. This algorithm could be deployed on top of any of the network protocols of prior work.
As an additional complication, replicas might adjust their reported orderings in response to strategic considerations. One desirable property of an order aggregation rule is that the influence of a (potentially colluding) subset of replicas is bounded and precisely quantified. Mirroring the prior work, we write here that it is “faulty” behavior for a replica to report anything other than the order in which it observes transactions arriving over the network.
There is a wide body of social choice literature on this underlying aggregation problem. Our goal with this work is to demonstrate an application for our streaming version of the problem, and to demonstrate the value of using social choice results in this application by targeting the precise desiderata raised in prior work. There are many other natural desiderata, notions of fairness, and aggregation rules that may be practically useful. The application of social-choice style aggregation rules in this streaming setting poses a number of interesting open questions. For example, if replicas have distinct financial motivations, or accept bribes from clients to order transactions in specific ways, is there an aggregation rule that maximizes (perhaps approximately) social welfare? For a reader coming from social choice theory, “transaction” could be replaced by “candidate”, and “an ordering vote” by “a ranking.”
1.1. Our Results
To the best of our knowledge, this work is the first to directly pose the question as a streaming version of the classic preference aggregation problem. We first observe that for an ordering method to be well-defined on a countably infinite set, it suffices for there to exist a “monotonic” and “asymptotically live” algorithm that implements the ordering method on finite, initial segments of the input. This algorithm must not undo a decision when its (finite) input is extended (”monotonicity”) and must eventually output every transaction (”asymptotic liveness”). The core challenge in building such an algorithm is in determining when an algorithm has enough information to make a decision that is consistent with its hypothetical output on any extension of its input.
Prior work of kelkar2020order defines a notion of “$\gamma$-batch-order-fairness” (reproduced here as Definition 4.1). The output ordering is divided into batches, and if $\gamma n$ replicas receive a transaction $\tau$ before another transaction $\tau^{\prime}$, then $\tau$ cannot be in a later batch than $\tau^{\prime}$. However, transactions within a batch can be ordered arbitrarily. Unfortunately, this definition is vacuously satisfiable by an arbitrarily large batch. We first strengthen this definition in §4 to precisely capture the intuition that batches should be minimal. Minimality must be defined carefully in the presence of faulty replicas; we show that $\gamma$-batch-order-fairness cannot be satisfied simultaneously with exact notions of batch minimality and faulty replicas.
We call our definition $(\gamma,\delta)$-minimal-batch-order-fairness (Definition 4.4). Simply put, if a $\frac{1}{2}<\gamma\leq 1$ fraction of replicas vote for $\tau$ before $\tau^{\prime}$ (henceforth; $\tau\prec\tau^{\prime}$) then the output ordering should include $\tau\prec\tau^{\prime}$ unless there is a sufficiently strong reason to put $\tau^{\prime}\prec\tau$—that is, a sequence of transactions $\left(\tau_{1},\tau_{2},\ldots,\tau_{k-1},\tau_{k}\right)$, with $\tau_{1}=\tau$ and $\tau_{k}=\tau^{\prime}$, where at least a $\gamma-2\delta$ fraction of replicas vote for $\tau_{i}\prec\tau_{i+1}$ for $11\leq i\leq k-$ 1. Motivating this definition is the fact that a $\delta$ fraction of faulty replicas can, at most, increase or decrease the fraction of replicas that report $\tau\prec\tau^{\prime}$ by $\delta$.
As an example, given $f$ faulty replicas out of $n$ total replicas and a fixed $\gamma$, the protocol of kelkar2020order satisfies $(\gamma,\frac{f}{n})$-minimal-batch-batch-order-fairness (Lemma 4.3). Unfortunately, that protocol relies on an explicit choice of a parameter $\gamma$ and an explicit bound on $f$, and it is not at all clear whether a higher or lower parameter $\gamma$ gives a stronger guarantee (§11 gives an example). By contrast, the algorithm we give here achieves $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness for every $\frac{1}{2}<\gamma\leq 1$ simultaneously, and for any number of faulty replicas. Fairness guarantees smoothly degrade as the number of faulty replicas increases.  [^1] Our algorithms do not require knowledge of a bound on the number of faulty replicas. This number may be separate from a bound on the number of faulty replicas in an underlying communication protocol. Standard Byzantine-fault tolerant communication protocols require $n\geq 3f+1$.
This work adapts the ordering known as Ranked Pairs (tideman1987independence) to our streaming setting. Informally, Ranked Pairs operates on a weighted, directed graph, where each vertex is a transaction and an edge from $\tau$ to $\tau^{\prime}$ has weight equal to the fraction of replicas that report $\tau\prec\tau^{\prime}$. Prior work (kelkar2020order; cachin2022quick), by contrast, operates on the unweighted, directed graph that is derived by taking the weighted graph which our work uses, removing all edges below weight $\gamma$, and dropping the weights on the remaining edges. This technical difference is the key to enabling our stronger results.
Analysis of this directed graph gives an important structural lemma about what pieces of information Ranked Pairs uses to determine its output. This lemma allows a streaming version of Ranked Pairs to carefully track how uncertainty about as-yet-unseen transactions propagates throughout Ranked Pairs’s algorithm. Ranked Pairs iterates over edges in order of edge weight. We find that the tiebreaking rule in this ordering is a critical part of the algorithm. Using a naive deterministic rule, our streaming algorithm might never output any transactions. However, we then show in §7 that deliberate construction of the tiebreaking rule guarantees asymptotic liveness.
Finally, a synchronous network assumption combined with our tiebreaking guarantees that our algorithm outputs each transaction after a finite amount of time. If at most $\Delta$ time elapses between when a transaction $\tau$ is sent and when every honest replica votes on it, for example, our algorithm outputs $\tau$ after a delay of at most $O(n\Delta)$. This can be reduced to $O(k\Delta)$, by rounding edge weights to the nearest $\frac{1}{k}$— although this rounding reduces the fairness guarantee to $(\gamma,\frac{f}{n}+\frac{1}{2k})$-minimal-batch-order-fairness (again for every $\frac{1}{2}<\gamma\leq 1$ simultaneously, and any $f$).
2. Preliminaries and System Model
We consider a model in which there are $n$ replicas cooperating to develop a total ordering of transactions. Transactions are received over the network from clients. We say that the ordering in which a replica receives transactions is that replica’s observed ordering.
Definition 2.1 (Ordering Preference).
An Ordering Preference on a (finite or countably-infinite) set of transactions is a total ordering 222 Isomorphic to a subset of $\omega$ $\hat{\sigma_{i}}=\left(\tau_{i_{1}}\prec\tau_{i_{2}}\prec\tau_{i_{3}}\prec\ldots\right)$
However, at any finite time, each replica can have received only a finite number of transactions. Replicas periodically submit these finite “ranking votes” to a ranking algorithm.
Definition 2.2 (Ordering Vote).
A replica $i$ submits to a ranking algorithm an ordering vote on a set of $k$ transactions, $\sigma=\left(\tau_{i_{1}},...\tau_{i_{k}}\right)$. (where $\tau_{i_{j_{1}}}\prec\tau_{i_{j_{2}}}$ for $j_{1}<j_{2}$).
We say that $\sigma^{\prime}$ extends $\sigma$ if $\sigma$ is an initial segment of $\sigma^{\prime}$ (as a convention, $\sigma$ extends itself).
Definition 2.3 (Ranking Algorithm).
A deterministic ranking algorithm $\mathcal{A}(\sigma_{1},\ldots,\sigma_{n})$ takes as input an ordering vote from each replica and outputs an ordering $\sigma$ on a subset of the transactions in its input.
The output need not include every transaction in the input.
We say that a replica is honest if its true, observed ordering always extends its ordering vote, and if whenever it submits a vote, it contains all transactions that the replica has observed at that time. Otherwise, the replica is faulty. We assume that at most $f$ of the $n$ replicas are faulty, although our algorithms do not require knowledge of $f$. 333 Of course, $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness (Definition 4.4) is only meaningful if $f<\frac{n}{2}$.
In the rest of this work, we assume that every replica eventually votes on every transaction. This may not necessarily hold for faulty replicas. A precise choice of response to this type of faulty behavior will depend on the network model and communication protocol, but, informally, if a replica fails to vote on a transaction within a “sufficient” time, it suffices for the honest replicas to fabricate a vote on behalf of the faulty replica. We address this difficulty in §8.
3. Related Work
3.1. Order-Fairness
Closest to this work are Aequitas (kelkar2020order) and Themis (themis), which propose the definition of $\gamma$-batch-order-fairness and instantiate ordering protocols based on aggregating ordering votes into a total order. Aequitas achieves $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness for a fixed choice of $\gamma$, but is not asymptotically live. §4.3 gives a more detailed comparison to these works.
Cachin et al. (cachin2022quick) gives a notion of “$\kappa$-differential-order-fairness;” Theorem D.1 of (themis) shows an equivalence between this definition and $\gamma$-batch-order-fairness, for a particular choice of $\gamma$ (as a function of $\kappa$).
Our definition of $(\gamma,\delta)$-minimal-batch-order-fairness (Definition 4.4) strengthens the (vacuously satisfiable) $\gamma$-batch-order-fairness definition used in prior work. §4 discusses the details of this relationship.
Concurrently with this work, vafadar2023condorcet observe that an adversary can influence the output ordering of many protocols (including that of kelkar2020order) by sending extra transactions (which can force batches to be combined together). This eliminates the constraints on the output that Definition 4.1 gives (using an example akin to Example 11.2). Lemma 4.3 identifies a similar problem. This observation is a corollary of the fact that the protocol of kelkar2020order (as well as the Ranked Pairs algorithm studied here) does not satisfy the Irrelevance of Independent Alternatives axiom (arrow1950difficulty).
vafadar2023condorcet also propose using Ranked Pairs to sort transactions within batches (as, e.g., those output by a protocol like Themis (themis)). However, no modification to the ordering of transactions within a batch output in Aequitas (kelkar2020order) can give the Ranked Pairs ordering, because Ranked Pairs might need to interleave batches that Aequitas considers incomparable (as demonstrated in §12). This strategy also cannot mitigate Aequitas’s lack of asymptotic liveness. This mitigation, if applied to Aequitas, is akin to dropping all edges of weight less than $\gamma$ before running Ranked Pairs. This strategy, therefore, would achieve $(\gamma^{\prime},\frac{f}{n})$-minimal-batch-order-fairness for all $\gamma^{\prime}\geq\gamma$ (a weaker guarantee than what our protocol provides).
As discussed in §4.3, the batches in Themis (themis) are defined slightly differently than in Definition 4.1. Themis, when computing a graph of ordering dependencies, adds a directed edge between every pair of vertices, choosing the direction of an edge by majority vote (ignoring the parameter $\gamma$, which is considered only on edges adjacent to transactions that some replicas have not yet reported observing). In so doing, it merges together any two batches unless, for any transaction $\tau$ in one and $\tau^{\prime}$ in the other, a majority of replicas vote $\tau\prec\tau^{\prime}$ (which bypasses the counterexample of §12).
Additionally, Themis’s (themis) round-based sequencing protocol presents a different problem insurmountable by any process like the one suggested by vafadar2023condorcet that sorts batches separately. Namely, Themis may finalize the composition of a batch before has enough information to do guarantee that, according to the Ranked Pairs ordering, all transactions in the batch must come before all those not in the batch (and not already in the output). This means that Themis may output two batches in sequence, such that running Ranked Pairs on each batch separately produces a different output than running Ranked Pairs on the union of the two batches together. We give a concrete example of this phenomenon in §13.
aequitaspermissionless instantiate Aequitas in a setting where the number of replicas is not known.
zhang2020byzantine adds timestamps to ordering votes and sorts transactions by median reported timestamp to provide a linearizability guarantee between transactions. This guarantee is incomparable to $(\gamma,\delta)$-minimal-batch-order-fairness.
3.2. Social Choice
Classical work of social choice studies the problem of choosing an order between a fixed, finite set of candidates given a complete set of preferences from each voter (arrow1950difficulty). Prior work studies related models in a limited-information setting. Lu and Boutilier (lu2013multi), Ackerman et. al (ackerman2013elections), and Cullinan et. al (cullinan2014borda) study the setting where a social choice rule has incomplete access (a partial ordering) to a voter’s preferences over a finite set of candidates. Conitzer and Sandholm (conitzer2005communication) study the communication complexity of a variety of voting rules. Fain et. al (fain2019random) study a randomized ordering rule that requires only a constant number of queries to voter preferences.
Fishburn (fishburn1970arrow) showed that Arrow’s impossibility theorem does not hold when the set of voters is infinite, although Kirman and Sondermann (kirman1972arrow) find “dictatorial sets” of voters. Grafe and Grafe (grafe1983arrow) extend these results to the case of infinitely many alternatives, subject to a continuity condition on the space of ranking preferences. Chichilnisky and Heal (chichilnisky1997social) and Efimov and Koshevoy (efimov1994topological) study the types of rules admissible in the infinite voters setting, given a topology on ordering preferences.
crisman2022voting study the preference aggregation problem where the output must be a cyclic permutation of the candidates.
3.3. Front-Running
The study of order fairness was initially motivated (as in, e.g., kelkar2020order) by the goal of limiting an adversary’s ability to order transactions (“front-running”) in public blockchains for their own financial benefit (daian2020flash). Additional approaches to this front-running problem include commiting to a transaction ordering before any replica can know transaction contents, using threshold encryption or commit-reveal schemes (malkhi2022maximal; clineclockwork; zhang2022flash). Li et. al (li2023transaction) study the problem of verifiably computing an ordering within an encrypted hardware enclave (costan2016intel). These could be applied on top of a streaming ordering algorithm.
Kavousi et. al (kavousi2023blindperm) randomly shuffle blocks of encrypted transactions. Some decentralized exchanges (ramseyer2023speedex; cowswapproblem; penumbraswap) process transactions in unordered batches, thereby eliminating the need to compute an ordering. ferreira2022credible construct an exchange-specific ordering that provides traders with certain worst-case price guarantees. mamageishvili2023buying use a median-timestamp mechanism akin to zhang2020byzantine, but allow clients to pay to reduce its reported timestamp.
4. Defining Fair Ordering
4.1. Batch Order Fairness
Our discussion here follows that of (kelkar2020order; themis; cachin2022quick), but gives, unlike prior work, a definition that is meaningful and achievable.
Honest replicas might receive transactions over the network in a different order. To resolve disagreements “fairly,” recent work (kelkar2020order) proposes a notion known as “$\gamma$-Batch-Order-Fairness,” which we reproduce below from Definition III.1 of themis (a more complicated but functionally similar definition is used in kelkar2020order).
Definition 4.1 ($\gamma$-batch-order-fairness).
Suppose that $\tau$ and $\tau^{\prime}$ are received by all nodes. If $\gamma n$ nodes received $\tau$ before $\tau^{\prime}$, then a ranking algorithm outputs $\tau$ no later than $\tau^{\prime}$.
Aequitas (kelkar2020order) outputs transactions in batches, and “no later than” means “in the same batch.”Transactions within a batch are ordered arbitrarily. Themis (themis) outputs a total ordering, with the assertion that this total ordering could be divided into contiguous (disjoint) batches satisfying this property.
One challenge is the parameter $\gamma$. It is not obvious if a higher or lower $\gamma$ gives stronger fairness properties. Indeed, there are simple examples (with no faulty replicas) where Definition 4.1 (actually the stronger Definition 4.4) only implies ordering restrictions at high values of $\gamma$, others at only low values, and still others at only intermediate values (§11). Prior work relies on a fixed choice of $\gamma$, but our algorithms satisfy Definition 4.4 for all values of $\gamma$ simultaneously.
4.2. Batch Minimality
Unfortunately, Definition 4.1 is insufficient to provide meaningful ordering guarantees. A protocol that outputs all transactions in the same batch vacuously satisfies this definition, and if transactions within a batch can be ordered arbitrarily, then every ordering satisfies Definition 4.1 for every $\gamma$.
Prior work (themis) recognizes the need for “minimality”.However, the notion of minimality in prior work is not achievable in the presence of faulty replicas (Theorem 4.2). Towards a feasible notion of minimality, we first define a graph denoting ordering dependencies.
Definition 4.2 (Ordering Graph).
Suppose that replicas report ordering votes $(\sigma_{1},...,\sigma_{n})$ on a set of transactions $V$. The Ordering Graph $G(\sigma_{1},...\sigma_{n})$ is a complete, weighted directed graph with vertex set $V$ and, for each transaction pair $(\tau,\tau^{\prime})$, an edge of weight $w(\tau,\tau^{\prime})=\alpha$ if $\alpha n$ replicas report receiving $\tau$ before $\tau^{\prime}$.
Any notion of “minimal” should at least require that when there are no cycles in ordering dependencies, the output is consistent with every ordering dependency (as in Theorem IV.1 part 1, Themis (themis)). Exact $\gamma$-minimality generalizes this notion to the case where cycles in ordering dependencies exist.
Definition 4.3 (Exact $\gamma$-minimality).
For any pair of transactions $\tau,\tau^{\prime}$, if $\gamma n$ replicas receive $\tau\prec\tau^{\prime}$ but $\tau^{\prime}$ is output before $\tau$, then there exists a sequence of transactions $\tau^{\prime}=\tau_{1},...,\tau_{k}=\tau$ where $\gamma n$ replicas receive $\tau_{i}\prec\tau_{i+1}$ and $\tau_{i}$ is output before $\tau_{i+1}$ for all $i$.
Unfortunately, this notion of minimality is impossible to achieve (for any $\gamma$) in the face of any faulty replicas. The condition on $f$ in Theorem 4.2 is a technical artifact of a construction in the proof (related to the pigeonhole principle). For some parameter settings, such as $n=7$ with $\gamma=\frac{2}{3}$, the condition admits $f=1$.
{theoremrep}
No protocol can achieve $\gamma$-batch-order-fairness with exactly $\gamma$-minimal output batches for any $n$ and $f$ greater than $n\mod(n-\lceil\gamma n\rceil+1)$ (for $\gamma>1/2$).
In fact, the construction in the proof contains no cycles in the ordering dependency graph of weight $\gamma$, thereby applying even to a weaker version of Definition 4.3 that applies only to the case where the ordering graph, restricted to edges of weight at least $\gamma$, has no cycles.
Proof.
Define m=$\lceil\gamma n\rceil$, and $k=\lfloor\frac{n}{n-m+1}\rfloor$. Consider the process of choosing an ordering between $k$ transactions $\tau_{1},\tau_{2},...,\tau_{k}$. Assume that every node submits a vote on transaction ordering (i.e. the theorem holds even when faulty nodes are required to submit a vote).
An ordering dependency arises, therefore, if and only if there are at least $m$ votes for some $\tau$ before another $\tau^{\prime}$.
Divide the set of nodes into disjoint groups $G_{i}$ for $1\leq i\leq k$ of size $n-m+1$. The remainder are the faulty nodes.
Suppose that nodes in group $i$ receive transactions in order $\tau_{i}\prec...\prec\tau_{k}\prec\tau_{1}\prec...\prec\tau_{i-1}$, and that the faulty nodes receive the same ordering of transactions as the nodes in group $G_{1}$.
As such, the transaction pairs $\tau_{i},\tau_{j}$ that receive at least $m$ votes are exactly those with $i<j$, and there are no cycles in this ordering dependency graph. Thus, any protocol must output the ordering $\tau_{1},...,\tau_{k}$.
However, for any protocol that lacks knowledge of which nodes are faulty, this scenario is indistinguishable from one where the faulty nodes had received transactions in the ordering received by group $G_{2}$, but reported the ordering observed in $G_{1}$. As such, the protocol would necessarily output the ordering observed by $G_{1}$.
However, in this case, the only correct output would have been $\tau_{2},...,\tau_{k},\tau_{1}$, as only the transaction pairs $\tau_{i},\tau_{j}$ with $i<j$ for $i\neq 1$ or $j=1$ received $m$ votes for $\tau_{i}\prec\tau_{j}$. ∎
The core difficulty is that faulty replicas can cause the weight on an edge observed by any algorithm to be different from the true, ground truth weight. Intuitively, if a protocol has a reliable communication layer (i.e. every honest replica is able to submit a vote on its ordering preference), then the worst that a faulty replica can do is to misreport its ordering preferences. If there are $f$ faulty replicas, then the faulty nodes can artificially reduce or increase the fraction of replicas that vote for $\tau\prec\tau^{\prime}$ by at most $\frac{f}{n}$. As a protocol cannot distinguish a faulty replica from an honest one by its votes, we therefore must take some error in edge weights into account in our definition of minimality.
Definition 4.4 ($(\gamma,\delta)$-minimal-batch-order-fairness).
An ordering is $(\gamma,\delta)$-minimally-batch-order-fair if, for any transaction pair $(\tau,\tau^{\prime})$ that is received in that order by at least $\gamma n$ replicas but output by the protocol in the reverse ordering, then there is a sequence of transactions $\tau^{\prime}=\tau_{1},...,\tau_{k}=\tau$ where at least $(\gamma-2\delta)n$ replicas receive $\tau_{i}\prec\tau_{i+1}$ and $\tau_{i}$ is output before $\tau_{i+1}$.
Definition 4.4 captures the notion that a protocol cannot distinguish between a $\delta$-fraction of replicas misreporting a transaction ordering. Given this indistinguishable $\delta$ fraction, the protocol outputs minimally-sized batches. Definition 4.3 corresponds to the case of $\delta=0$.
This definition does not explicitly discuss “batches” or “minimality,” but approximately minimal batches (the strongly connected components of Lemma 4.2) can be recovered from any ordering satisfying it. The second condition of Lemma 4.2 limits the size of output batches.
{lemmarep}
Suppose that an output ordering satisfies $(\gamma,\delta)$-minimal-batch-order-fairness. Compute the ordering graph $\hat{G}$, drop all edges with weight below $\gamma-\delta$, and compute the strongly connected components of the remainder.
  • If $\gamma n$ replicas received $\tau\prec\tau^{\prime}$, then either $\tau$ and $\tau^{\prime}$ are in the same strongly connected component, or all transactions in the component containing $\tau$ are output before any transactions in the component containing $\tau^{\prime}$.
  • If $\gamma n$ replicas receive $\tau\prec\tau^{\prime}$ and there is no sequence of transactions $\tau^{\prime}=\tau_{1},...,\tau_{k}=\tau$ where at least $(\gamma-2\delta)n$ replicas receive $\tau_{i}\prec\tau_{i+1}$, then all transactions in the component containing $\tau$ are output before any transactions in the component containing $\tau^{\prime}$.
Proof.
If at least $\gamma n$ replicas receive $\tau\prec\tau^{\prime}$, then the edge $(\tau,\tau^{\prime})$ is included in the thresholded ordering graph, and if less than $(\gamma-2\delta)$ replicas receive $\tau\prec\tau^{\prime}$, then the edge is not included.
Thus, if at least $\gamma n$ replicas receive $\tau\prec\tau^{\prime}$, then either $\tau$ and $\tau^{\prime}$ are in the same strongly connected component, or there is an edge from the component containing $\tau$ to that containing $\tau^{\prime}$. And if, additionally, there is no sequence of transactions from $\tau^{\prime}$ to $\tau$ as in the lemma statement, then $\tau$ and $\tau^{\prime}$ must be in different components.
If there is an edge from one strongly connected component to another, then all transactions in the first component must be output before any in the second (or else a violation of Definition 4.4) would occur. ∎
There is one subtle difference between this definition and (an analogue of) the notion of “minimality” in Theorem IV.1 of Themis, and the sequence of batches output in Aequitas. When there are two disjoint strongly connected components with no dependencies (of strength at least $\gamma$) from one to the other, Definition 4.4 allows the output to interleave these components. This may be strictly required (§12).
4.3. Comparison to Prior Work
Before constructing our protocol, we first apply Definition 4.4 to related prior work.
{lemmarep}
Aequitas with parameters $(\gamma,f)$ achieves $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness.
Note that Aequitas requires $\gamma-\frac{1}{2}>\frac{2f}{n}$. Lemma 4.3 is a stronger version of Theorem 6.6 of (kelkar2020order).
Proof.
Aequitas considers a transaction pair $(\tau,\tau^{\prime})$ as an ordering dependency if at least $\gamma n-f$ report receiving $\tau$ before $\tau^{\prime}$ (part 3.b.ii of §6.2 of (kelkar2020order)). This ensures that the protocol considers every ordering dependency that truly receives at least $\gamma n$ votes, and might include any edge that receives at least $\gamma n-2f$ votes.
Aequitas sequentially outputs minimal cycles in its computed graph of ordering dependencies (not Definition 4.2), ordering within each cycle arbitrarily. This means that if some pair that receives at least $\gamma n$ votes is output in reversed order, then it must be part of a cycle of dependencies, and each of these dependencies must have received at least $\gamma n-2f$ votes in order to be included in Aequitas’s computed graph. ∎
By contrast, Ranked Pairs (Theorem 5) achieves $(\gamma,\frac{f}{n})$-minimal-batch-fairness for every $\gamma$ simultaneously, and for any $f$ (our work does not require knowledge of a bound on $f$). Aequitas also is not asymptotically live (Definition 6.2).
Themis claims (Theorem IV.1, (themis)) to output a sequence that can be partitioned into minimal batches, which appears to contradict Theorem 4.2. However, Themis relaxes its notion of an “ordering dependency”.
Where $\gamma$-batch-order-fairness would create a dependency for $\tau$ not after $\tau^{\prime}$ if at least $\gamma n$ replicas receive $\tau$ before $\tau^{\prime}$, the relaxation implied by Definition III.1 of (themis) creates such a dependency if at least $n(1-\gamma)+1$ receive $\tau$ before $\tau^{\prime}$ (that is, it cannot be the case that more than $\gamma n$ receive $\tau^{\prime}$ before $\tau$). Under this definition, the construction in Theorem 4.2 constitutes what the authors of (themis) call a “weak Concorcet Cycle”.
Extra ordering dependencies, however, can artificially merge batches together, giving counterintuitive results. {lemmarep} In Themis (themis), even if all nodes report $\tau$ before $\tau^{\prime}$ and there are no faulty replicas nor cycles in ordering dependencies (as defined in Definition 4.1) for any $\gamma>1/2$, the output may put $\tau^{\prime}$ before $\tau$.
Lemma 4.3 follows from the fact that when two transactions $\tau,\tau^{\prime}$ receive $n-f$ votes in one round of Themis, it determines whether to consider as a dependency $\tau\prec\tau^{\prime}$ or $\tau^{\prime}\prec\tau$ by majority vote (disregarding $\gamma$). The minimality statement of Theorem IV.1 of (themis) is with regard to this weaker notion of dependency.
Proof.
Suppose that the number of nodes is even, and break the nodes into two groups. Group one receives transactions $\tau_{1}\prec\tau_{2}\prec\tau_{3}$, while group two receives transactions $\tau_{3}\prec\tau_{1}\prec\tau_{2}$.
All nodes submit their local orderings to the protocol correctly. Themis proceeds in rounds; the round under consideration consists only of these three transactions.
Themis builds a graph of what it considers ordering dependencies on the transactions in a round (Figure 1, part 1, (themis)). Let $w(\tau,\tau^{\prime})$ be the number of replicas that vote for $\tau$ before $\tau^{\prime}$.
An ordering dependency between $\tau$ and $\tau^{\prime}$ is included if (1) $w(\tau,\tau^{\prime})\geq w(\tau^{\prime},\tau)$ and (2) $w(\tau,\tau^{\prime})\geq n(1-\gamma)+f+1$. Themis assumes $n>4f/(2\gamma-1)$, or equivalently, $2f<n(\gamma-1/2)$, so $(1-\gamma)+(f+1)/n\leq(1-\gamma)+(2f)/n\leq(1-\gamma)+(\gamma-1/2)\leq 1/2$. Ties are broken in an unspecified (deterministic) manner.
As such, (depending on the tiebreaking), Themis may consider as valid ordering dependencies the pairs $(\tau_{1},\tau_{2})$, $(\tau_{2},\tau_{3})$, and $(\tau_{3},\tau_{1})$. Within a strongly connected component of its dependency graph, Themis computes a Hamiltonian cycle, then outputs transactions by walking arbitrarily along that cycle. As such, a possible output of Themis is the ordering $\tau_{2}\prec\tau_{3}\prec\tau_{1}$.
Thus, all nodes received $\tau_{1}$ before $\tau_{2}$, and there were no cycles of ordering dependencies for any $\gamma>1/2$, but the protocol could output $\tau_{2}$ before $\tau_{1}$.
Note that Themis only waits for $n-f$ votes from replicas before proceeding, not the full $n$. This difference is immaterial if $f=0$. However, the construction above works only if the numbers of votes for $\tau_{3}\prec\tau_{1}$ and for $\tau_{2}\prec\tau_{3}$ are both at least $n(1-\gamma)+(f+1)$. Tighter analysis shows that $n(1-\gamma)+(f+1)=n(1-\gamma)+(2f)-(f-1)\leq n(1-\gamma)+n(\gamma-1/2)-(f-1)=n/2-(f-1)<(n-f)/2$, so the construction still works.
For $\gamma$ larger than $1/2+\varepsilon$ for some small $\varepsilon$ and sufficiently large $n$, this construction could be repeated with more transactions in the cycle. This eliminates the need to abuse a tiebreaking rule when votes are evenly divided. ∎
5. Ranked Pairs Voting
We turn now to the Ranked Pairs method (tideman1987independence). Ties in edge weights are broken according to an arbitrary (deterministic) rule.
Input: An ordering vote from each replica $\{\sigma_{i}\}_{i\in[n]}$
$G=(V,E,w)\leftarrow G(\sigma_{1},...,\sigma_{n})$ ;
  /* Compute ordering graph */
$H\leftarrow(V,\emptyset)$ ;
  /* Initialize empty graph on same vertex set */
Sort $E$ by edge weight
foreach $e\in E$ do
       Add $e$ to $H$ if it would not create a directed cycle in $H$
end foreach
Output: The topological sort of $H$
  /* The loop considered every edge, so the topological sort of $H$ is unique */
ALGORITHM 1 Ranked Pairs Voting
{theoremrep}
Given a ordering vote (on every transaction in a finite set) from every replica, Ranked Pairs Voting simultaneously satisfies $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness for every $\gamma$, and does not depend on any fixed bound on $f$.
Lemma 4.2 and Theorem 5 together imply an important property. The output of Ranked Pairs can be divided into batches consistent with $\gamma$-batch-order-fairness for any $\gamma$, not just a single $\gamma$ (subject to the interleaving discussed in §4.2). Unlike prior work, there is no need to choose a $\gamma$ or order transactions within a batch.
Proof.
Ranked Pairs Voting ensures that if some edge of weight $\gamma$ is not included in the output graph $G$ — that is to say, if $\gamma n$ nodes vote for $\tau$ before $\tau^{\prime}$, but the output ordering has $\tau^{\prime}\prec\tau$ — then there must be a directed path of edges in $G$ already from $\tau^{\prime}$ to $\tau$. As the algorithm looked at these edges before the current edge, these edges must have weight (as observed by the algorithm) at least $\gamma$.
Note that if there are $f$ faulty nodes, these nodes can, at most, adjust the weight (observed by the algorithm) on any particular transaction pair by at most $\frac{f}{n}$.
Thus, an edge with true weight $\gamma$ is only reversed if there exists a path in the ordering graph (that is included in $G$) in the opposite direction of minimum true weight at least $\gamma-\frac{2f}{n}$.
Nowhere in the algorithm do its choices depend on $\gamma$ or $f$ (or $n$). ∎
6. Streaming Ranked Pairs
We now turn to the streaming setting that is the focus of our work. There may be an infinite set of transactions to consider overall, but at any finite time, an algorithm must compute an ordering on a subset of the transactions that have been seen thus far. In a real-world system, such an algorithm would be run with some periodic frequency, and replicas would periodically append to their ordering votes.
Our algorithms must be monotonic (Definition 6.1)—if replicas extend their ordering votes, the algorithm, when run again on the extensions, can only extend its prior output.
Definition 6.1 (Monotonicity).
A sequencing algorithm $\mathcal{A}(\cdot,\ldots,\cdot)$ is monotonic if, given two sets of ordering votes $\left(\sigma_{1},\ldots,\sigma_{n}\right)$ and $\left(\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{n}\right)$, such that $\sigma^{\prime}_{i}$ extends $\sigma_{i}$ for all $i\in[n]$, $\mathcal{A}(\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{n})$ extends $\mathcal{A}(\sigma_{1},\ldots,\sigma_{n})$.
Observe that if a sequencing algorithm is monotonic, it implies a well-formed definition for aggregating a set of orderings on countably infinite sets of transactions, not just on finite inputs. Specifically, $\tau$ comes before $\tau^{\prime}$ in the overall (infinite) case if there exists a finite subset of the input orderings such that the algorithm puts $\tau\prec\tau^{\prime}$ in its output.
A non-vacuous condition is also required—the algorithm that never outputs anything is monotonic.
Definition 6.2 (Asymptotic Liveness).
A sequencing algorithm $\mathcal{A}(\cdot,\ldots,\cdot)$ is asymptotically live if, given any set of countably infinite ordering votes $(\hat{\sigma}_{1},\dots,\hat{\sigma_{n}})$ and any transaction $\tau$ in those votes, there exists an $N$ such that when each $\hat{\sigma_{i}}$ is trunctated to the first $N$ elements of the ordering to produce $\sigma_{i}$, $\tau$ is included in $\mathcal{A}(\sigma_{1},\ldots,\sigma_{n})$.
For expository simplicity, we present our algorithm in two steps. First, we give a streaming algorithm (Algorithm 2) that is monotonic but not asymptotically live. Then, we give a modification that ensures asymptotic liveness (Algorithm 3).
6.1. Decisions Are Local
Implementing Ranked Pairs voting in this streaming setting requires first understanding when the algorithm chooses or rejects an edge in the ordering graph.
{lemmarep}
If Ranked Pairs rejects an edge $(\tau,\tau^{\prime})$ of weight $\alpha$, then it must choose some directed path of edges from $\tau^{\prime}$ to $\tau$, where each edge has weight at least $\alpha$.
Proof.
If Ranked Pairs rejects an edge $(\tau,\tau^{\prime})$ of weight $\alpha$, then adding the edge would have created a cycle among the edges the algorithm had already chosen. Ranked Pairs looks at edges in order of weight, so all of the edges on the already chosen path from $\tau^{\prime}$ to $\tau$ must have weight at least $\alpha$. ∎
{observationrep}
Ranked Pairs chooses every edge of weight $1$.
Proof.
There cannot be any cycles of edges, all of weight $1$. Such a situation would require every replica to submit a cycle of preferences, which is impossible, given that each ordering vote $\sigma$ is a total, linear order. ∎
Furthermore, the set of vertices that such a path can visit are bounded to a local neighborhood of $\tau_{A}$ and $\tau_{B}$.
Definition 6.3 ().
Let $\tau$ be any transaction.
  1. (1)
    Let $P_{\tau}$ be the set of all preceeding transactions; that is, all $\tau^{\prime}$ with $w(\tau^{\prime},\tau)=1$.
  2. (2)
    Let $Q_{\tau}$ be the set of all concurrent transactions; that is, all $\tau^{\prime}$ with $10<w(\tau^{\prime},\tau)<$ 1.
  3. (3)
    Let $R_{\tau}$ be the set of all subsequent transactions; that is, all $\tau^{\prime}$ with $w(\tau^{\prime},\tau)=0$.
{lemmarep}
If ranked pairs chooses all edges in a path from $\tau_{B}$ to $\tau_{A}$, no transactions on the path are in $R_{\tau_{A}}$ or $P_{\tau_{B}}$.
Proof.
If the path visits a transaction $\tau$ in $R_{\tau_{A}}$, then it creates a cycle with the edge from $\tau_{A}$ to $\tau$ (which ranked pairs must choose, by Observation 6.1). If the path visits a transaction $\tau$ in $P_{\tau_{B}}$, then it creates a cycle with the edge from $\tau$ to $\tau_{B}$. ∎
Lemma 6.1 lets the streaming algorithm decide whether to choose an edge given only a finite amount of information. Once the algorithm sees a vote for a transaction $\tau$ from every replica, it can compute the weight in the ordering graph for every edge $(\tau,\tau^{\prime})$ and $(\tau^{\prime},\tau)$. Unseen transactions must be in $R_{\tau}$.
Even so, a streaming algorithm cannot always know whether or not Ranked Pairs would choose an edge. As such, we allow our algorithms to leave an edge in an “indeterminate” state, and design our algorithms to account for indeterminate edges when considering subsequent edges.
First, we construct an ordering graph that captures the information available to an algorithm at a finite time.
Definition 6.4 (Streamed Ordering Graph).
Suppose that each replica submits an ordering vote $(\sigma_{1},...,\sigma_{n})$. Let $V$ be the set of all transactions that appear in each vote, and let $\hat{v}$ be the “future” vertex.
The Streamed Ordering Graph is a weighted, directed, complete graph $\hat{G}=(V^{\prime},E)$ on vertex set $V\cup\{\hat{v}\}$.
For each $\tau$ and $\tau^{\prime}$, set weights $w(\tau,\tau^{\prime})$ and $w(\tau^{\prime},\tau)$ as in Definition 4.2. Set $w(\tau,\hat{v})=1$ for all $\tau$. If there exists $\tau^{\prime}$ that appears in the votes of some but not all replicas and which preceeds $\tau$ in at least one replica’s vote, set $w(\hat{v},\tau)=1$, and otherwise $w(\hat{v},\tau)=0$.
Conceptually, the “future” vertex $\hat{v}$ represents all transactions that have not received votes from all replicas. Implicitly, all edge weights not computable from the available information are upper-bounded by $1$. An implementation might compute a tighter upper bound on the weights of edges $(\hat{v},\tau)$. Subsequent arguments require only that the assigned weight is an upper bound of the true weight.
6.2. A Streaming Algorithm
We can now construct an algorithm to compute Ranked Pairs in our streaming setting. Instead of either including or rejecting every edge, Algorithm 2 has the option to declare an edge as indeterminate; that is, the algorithm is not able to determine whether to include or exclude the edge with the information in its input.
Input: An ordering vote from each replica $\{\sigma_{i}\}_{i\in[n]}$
$\hat{G}=(V,E,w)\leftarrow\hat{G}(\sigma_{1},...,\sigma_{n})$ ;
  /* Compute streamed ordering graph */
$H\leftarrow(V,\emptyset)$ ;
  /* Initialize empty graph on same vertex set */
foreach edge $(\tau,\hat{v})$ and $(\hat{v},\tau)$ with $w(e)=1$ do
       Add $e$ to $H$, and mark it as indeterminate
end foreach
Sort $E$ by edge weight, with a fixed tiebreaking rule.
foreach $e=(\tau_{i},\tau_{j})\in E$ do
       $U_{\tau_{i},\tau_{j}}\leftarrow(V\setminus(R_{\tau_{i}}\cup P_{\tau_{j}}))\cup\{\hat{v}\}$
        /* Recall that $\hat{v}$ is the ‘‘future’’ vertex of the streamed ordering graph */
       If there is a directed path in $H$ restricted to $U_{\tau_{i},\tau_{j}}$ from $\tau_{j}$ to $\tau_{i}$ where every edge is determinate, then do not include the edge in $H$.
       If there is no directed path in $H$ restricted to $U_{\tau_{i},\tau_{j}}$ from $\tau_{j}$ to $\tau_{i}$ where every edge is either determinate or indeterminate, add $(\tau_{i},\tau_{j})$ to $H$ and mark it as determinate.
       Otherwise, add $(\tau_{i},\tau_{j})$ to $H$ and mark it as indeterminate.
end foreach
Output: The topological sort of $H$ (with sorting comparisons restricted to determinate edges), up to but not including the first transaction $\tau$ bordering an indeterminate edge
ALGORITHM 2 Streaming Ranked Pairs
To prove correctness of this algorithm, it suffices to show that the output of the algorithm is monotonic and that the output is consistent with the non-streaming version of Ranked Pairs. By consistency, we specifically mean that at any point, if clients were to stop sending new transactions and all replicas eventually voted on every replica, Algorithms 1 and 2 would produce the same output.
{lemmarep}
Consider a counterfactual scenario where clients stop sending transactions, all replicas eventually receive every transaction, and every replica includes every transaction in its output vote.
Whenever Algorithm 2 includes a determinate edge in $H$ (resp. excludes an edge), that edge is included (resp. excluded) in the output of the non-streaming Ranked Pairs on the counterfactual input.
Proof.
Suppose that the lemma statement holds for all edges earlier in the Ranked Pairs ordering. Specifically, assume that Ranked Pairs would choose every higher determinate edge, not choose any higher rejected edge, and might or might not choose any higher indeterminate edge (and that all edges of higher weight are all either already rejected, chosen as determinate, or chosen as indeterminate).
By Lemma 6.1, Ranked Pairs would choose $(\tau_{i},\tau_{j})$ if and only if there does not exist a path of already chosen edges higher in the ordering from $\tau_{j}$ to $\tau_{i}$ that does not enter $R_{\tau_{i}}$ or in $P_{\tau_{j}}$.
If there exists a path of determinate edges in $U_{\tau_{i},\tau_{j}}$, then Streaming Ranked Pairs rejects $(\tau_{i},\tau_{j})$, and if there does not exist a path of determinate or indeterminate edges, then Streaming Ranked Pairs accepts $(\tau_{i},\tau_{j})$. Otherwise, the edge is left indeterminate.
Note that Streaming Ranked Pairs considers edges in the same ordering that Ranked Pairs would (relative to the restricted set considered). However, at any point, due to the initialization of $H$ with indeterminate edges, the set of edges in which the algorithm searches for a cycle might include more edges than those that would be considered by Ranked Pairs. However, these extra edges are indeterminate and can therefore only make Streaming Ranked Pairs choose indeterminate for a considered edge, satisfying the induction hypothesis.
Importantly, the weights of these extra edges upper bound their true (unknown) weights. And any cycle that would go through (an) as-yet-unseen transaction(s) maps to one (using indeterminate edges) through the future vertex $v$. As such, the set of paths considered in Streaming Ranked Pairs is always a superset of that considered by Ranked Pairs, and the difference between these sets is always made up of indeterminate edges.
As such, when the streaming algorithm does not choose to leave an edge as indeterminate, the streaming algorithm makes the same decisions as the non-streaming algorithm. Thus, the induction hypothesis holds for $(\tau_{i},\tau_{j})$.
The induction hypothesis clearly holds for the first edge considered, so the lemma holds. ∎
The same argument shows that Algorithm 2 is monotonic.
{corollaryrep}
Algorithm 2 is monotonic.
Proof.
The proof of Lemma 6.2 actually shows that an edge is marked determinate or rejected only if Ranked Pairs would make that decision on the edge on any counterfactual scenario that extends the input to the algorithm.
Lemma 6.2 directly implies that the output of Algorithm 2 matches that of Ranked Pairs.
{theoremrep}
The output of Algorithm 2 matches an initial segment of the ordering output by the non-streaming Ranked Pairs (on the counterfactual of Lemma 6.2).
Proof.
Note that every transaction bordering on an indeterminate edge must come strictly after (in the topological ordering) every transaction in the output of the algorithm.
By Lemma 6.2, then regardless of whether or not the indeterminate edges are chosen in Ranked Pairs, every transaction bordering an indeterminate edge must come after (in the true output of Ranked Pairs) all of the transactions output by the algorithm.
Within the set of output transactions, again because of Lemma 6.2, transactions must be ordered according to the true output of Ranked Pairs. ∎
Ranked Pairs has the property that, for any set of inputs $(\sigma_{1},\ldots,\sigma_{n})$, if it is run on a restriction of the inputs to the transactions that appear as an initial segment of its output on $(\sigma_{1},\ldots,\sigma_{n})$, its output is unmodified on the restricted inputs (for completeness, we give a proof in §10). As Algorithm 2 is monotonic, its output exactly matches the Ranked Pairs ordering when the input votes are restricted in this way.
7. Liveness and Efficiency
Input: An ordering vote from each replica $\{\sigma_{i}\}_{i\in[n]}$
Input: $\hat{H}$, the graph computed in the preceeding invocation of Algorithm 3
$\hat{G}=(V,E,w)\leftarrow\hat{G}(\sigma_{1},...,\sigma_{n})$ ;
  /* Compute streamed ordering graph */
$H\leftarrow(V,\emptyset)$ ;
  /* Initialize empty graph on same vertex set */
foreach edge $(\tau,\hat{v})$ and $(\hat{v},\tau)$ with $w(e)=1$ do
       Add $e$ to $H$, and mark it as indeterminate
end foreach
Add all determinate edges of $\hat{H}$ to $H$, marked determinate
foreach $\gamma$ in $(1,\frac{n-1}{n},\frac{n-2}{n},\ldots,0)$ do
       $E_{\gamma}\leftarrow\{e\in E~{}|w(e)=\gamma\}$, ordered arbitrarily
       repeat
             $e=(\tau_{i},\tau_{j})\leftarrow$ first element of $E_{\gamma}$
             $U_{\tau_{i},\tau_{j}}\leftarrow(V\setminus(R_{\tau_{i}}\cup P_{\tau_{j}}))\cup\{\hat{v}\}$
             If there is a directed path in $H$ restricted to $U_{\tau_{i},\tau_{j}}$ from $\tau_{j}$ to $\tau_{i}$ where every edge is determinate, then do not include the edge in $H$. Remove the edge from $E_{\gamma}$.
             Else if there is no directed path in $H$ restricted to $U_{\tau_{i},\tau_{j}}$ from $\tau_{j}$ to $\tau_{i}$ where every edge is either determinate or indeterminate, add $(\tau_{i},\tau_{j})$ to $H$ and mark it as determinate. Remove the edge from $E_{\gamma}$.
             Otherwise, defer $(\tau_{i},\tau_{j})$ to the end of the ordering on $E_{\gamma}$.
      until $E_{\gamma}$ is empty or no progress is made in one full loop over $E_{\gamma}$;
      Mark all remaining edges in $E_{\gamma}$ as indeterminate
end foreach
Output: The topological sort of $H$ (with sorting comparisons restricted to determinate edges), up to but not including the first transaction bordering an indeterminate edge
ALGORITHM 3 Streaming Ranked Pairs, Version 2
7.1. Efficient Marginal Runtime
Instantiating this protocol means running a ranking algorithm repeatedly (as replicas report additional information). Rather than recompute the streamed ordering graph and the status of every edge in each invocation, Algorithm 2 could use the output of a past invocation, restricted to the decisions that were determinate, as an oracle. Corollary 6.2 implies that determinate choices are consistent with any extension of the input, so this oracle does not change the output. However, the runtime now depends only on the number of new edges in the streamed ordering graph (§15).
7.2. Protocol Liveness
Algorithm 2 is not asymptotically live. §14 gives an instance on which the algorithm would never output any transaction.
That construction uses an adversarial tiebreaking rule, but Ranked Pairs requires only some tiebreaking rule. Our algorithm can therefore generate a tiebreaking rule dynamically. Specifically, when the algorithm visits edges of weight $\gamma=\frac{k}{n}$, it defers edges that would become indeterminate to the end of the ordering among edges of that weight.
Algorithm 3 constructs a (partial) ordering between edges that breaks ties between edges of the same weight. The output of this algorithm is, by the same arguments as in §6.2, consistent with the output of non-streaming Ranked Pairs, when using an tiebreaking rule consistent with this edge ordering. The oracle mechanism of §7.1 passes this tiebreaking information from one run of Algorithm 3 to the next.
{lemmarep}
Whenever Algorithm 3 includes a determinate edge in $H$, or excludes an edge from $H$, the non-streaming Ranked Pairs, when using the implied ordering, would make the same decision.
Proof.
The argument proceeds as in Lemma 6.2. Algorithm 3 induces a tiebreaking rule between edges of the same weight. If Algorithm 2 were given this tiebreaking rule, it would visit edges in the order in which Algorithm 3 makes decisions on edges (i.e., skipping the instances where Algorithm 3 defers an edge until later). As such, its output is consistent with the non-streaming Ranked Pairs using this implied ordering (by Theorem 6.2). ∎
Crucially, this small change enables the following structural lemma. We start with a useful definition.
Definition 7.1 ().
An edge $(\hat{\tau},\hat{\tau^{\prime}})$ is contemporary to an edge $(\tau,\tau^{\prime})$ if and only if $\hat{\tau}$ and $\hat{\tau^{\prime}}$ are contained in $U_{\tau,\tau^{\prime}}$.
{lemmarep}
If Algorithm 3 marks an edge $(\tau_{i},\tau_{j})$ with weight $\frac{k}{n}$ as indeterminate, there must be a path from $\tau_{j}$ to $\tau_{i}$ contained in $U_{\tau_{i},\tau_{j}}$ that contains an indeterminate edge of weight at least $\frac{k+1}{n}$.
By contrast, the best bound in Algorithm 2 is that there exists a contemporary indeterminate edge of weight $\frac{k}{n}$.
Proof.
By construction of the algorithm, if an edge $(\tau_{i},\tau_{j})$ is marked as indeterminate, then there must be a path in $U_{\tau_{i},\tau_{j}}$ already chosen (when this edge is visited) of determinate and indeterminate edges of weight at least $\frac{k}{n}$. Furthermore, any indeterminate edge must be of weight strictly more than $\frac{k}{n}$. When the algorithm looks for such a path, it has not yet marked any edges of weight $\frac{k}{n}$ as indeterminate, and if no such path exists, then the edge would not be marked as indeterminate. ∎
Lemma 7.2 implies that for any indeterminate edge, there is sequence of contemporary indeterminate edges of strictly increasing weights that causes the edge to be indeterminate.
{lemmarep}
Suppose Algorithm 3 marks an edge $e_{1}$ of weight $\gamma$ as indeterminate. Then there exists a sequence of indeterminate edges $e_{1},\ldots e_{k}$ of strictly increasing weight such that for each pair of edges $e_{i}=(\tau_{i},\tau_{j})$ and $e_{i+1}=(\tau^{\prime}_{i},\tau^{\prime}_{j})$ with $e_{i+1}$ contemporary to $e_{i}$, $\tau^{\prime}_{i}$ and $\tau^{\prime}_{j}$ are present in $U_{\tau_{i},\tau_{j}}$, and the source of $e_{k}$ is $\hat{v}$.
Proof.
Follows from repeated application of Lemma 7.2.
The algorithm starts with only edges leaving the “future” vertex $\hat{v}$ marked as indeterminate. These initial edges are the only indeterminate edges with weight $1$, and thus are the only ones that can form the root of any chain. ∎
None of these chains can be longer than $n$ edges, as each step increases the weight by at least $\frac{1}{n}$. This bound implies that Algorithm 3 eventually outputs every transaction.
{theoremrep}
Algorithm 3 is asymptotically live.
Proof.
Let $(\hat{\sigma_{1}},\dots,\hat{\sigma_{n}})$ be any set of (countably infinite) ordering preferences, and let $\tau$ be any transaction in those orderings.
Consider the set of all chains of edges $(e_{1},\ldots,e_{k})$ where $e_{i+1}$ is contemporary to $e_{i}$ (and of higher weight) and $e_{1}$ is adjacent to $\tau$, and $k\leq n$. Because each $U_{\tau,\tau^{\prime}}$ is finite, the number of such chains must be finite, and the number of transactions that appear in any of these chains must be finite.
On any input to Algorithm 3 that is a finite truncation of $(\hat{\sigma_{1}},\dots,\hat{\sigma_{n}})$, if an edge adjacent to $\tau$ is left indeterminate, then Lemma 7.2 implies that there must be a chain of indeterminate edges on this input with the first adjacent to $\tau$, the last adjacent to $\hat{v}$, and each edge contemporary to the previous. This chain, with the last edge dropped, must be one of the chains considered above. As such, there are only a finite number of transactions that could be in the last edge adjacent to $\hat{v}$. For each such transaction $\tau^{\prime}$, there can only be an edge from $\hat{v}$ to $\tau^{\prime}$ in the streamed ordering graph if there is some other transaction in the input that is ahead of $\tau^{\prime}$ in at least one replica’s vote. There can only be a finite number of these transactions.
Taking a union over a finite number of finite sets gives a finite set of transactions. There must therefore be some bound $M_{\tau}$ such that all of these transactions appear in every replica’s vote if $(\hat{\sigma_{1}},\dots,\hat{\sigma_{n}})$ is truncated to (at least) the first $M_{\tau}$ elements.
A transaction $\tau$ is not output by Algorithm 3 if the topological sort in the last step puts a transaction $\tau^{\prime}$ adjacent to an indeterminate edge ahead of $\tau$. But there are only a finite number of transactions that might ever be ahead of $\tau$ (specifically, $P_{\tau}$ and $Q_{\tau}$).
Let $N_{\tau}=\max_{\tau^{\prime}\in P_{\tau}\cup Q_{\tau}}M_{\tau^{\prime}}$. Then if $(\hat{\sigma_{1}},\dots,\hat{\sigma_{n}})$ is truncated to (at least) the first $N_{\tau}$, then $\tau$ must appear in the output of Algorithm 3. ∎
Additionally, under network synchrony, Lemma 7.2 bounds the temporal delay between when a transaction is sent by a client and when Algorithm 3 adds it to its output.
Assumption 7.2 (Synchronous Network).
If a client sends a transaction $\tau$ at time $t$, all replicas include $\tau$ in their ordering votes before time $t+\Delta$.
As such, replicas can disagree on the ordering between $\tau$ and $\tau^{\prime}$ only if they were sent at roughly the same time.
{observationrep}
Consider any two transactions $\tau_{i}$ and $\tau_{j}$ sent at times $t_{i}$ and $t_{j}$, respectively.
If $t_{i}+\Delta<t_{j}$, then the weight of the edge $(\tau_{i},\tau_{j})$ is $1$.
Proof.
If $t_{i}+\Delta<t_{j}$, then every replica receives and commits to a vote on $\tau_{i}$ before $\tau_{j}$ is sent, so $\tau_{j}$ must come after $\tau_{i}$ in every replica’s vote. ∎
{lemmarep}
Consider any three transactions $\tau$, $\tau_{i}$, and $\tau_{j}$ received by all replicas, that were sent by clients at times $t$, $t_{i}$, and $t_{j}$, respectively. Suppose $\tau\in U_{\tau_{i},\tau_{j}}$, and that the weight of $(\tau_{i},\tau_{j})$ is not 00.
Then $t_{j}-\Delta\leq t\leq t_{j}+2\Delta$.
Proof.
By construction, $\tau\notin R_{\tau_{i}}$, so $t\leq t_{i}+\Delta$. Furthermore, $\tau\notin P_{\tau_{j}}$, so $t+\Delta\geq t_{j}$. Furthermore, by Observation 7.2, $t_{j}+\Delta\geq t_{i}$.
Thus, $t_{j}-\Delta\leq t\leq t_{i}+\Delta\leq t_{j}+2\Delta$. ∎
This lemma and Lemma 7.2 give an overall time bound.
{theoremrep}
A transaction $\tau$ is contained in the output of the algorithm of Algorithm 3 after at most $(n+1)\Delta$ time.
Proof.
Let the current time be $T$, and let $\tau$ be sent at time $t$. Without loss of generality, assume $\tau$ has been voted on by every replica.
If an indeterminate edge exists to $\tau$, then there exists some transaction $\tau^{\prime}$ (sent at $t^{\prime}$) that has not been voted on by every replica, but which preceeds $\tau$ in some replicas’ votes. As such, $T-\Delta\leq t^{\prime}$. Then, as $\tau$ does not fully preceed $\tau^{\prime}$, it must be the case that $t^{\prime}-\Delta\leq t$ (so $T-2\Delta\leq t)$.
Lemma 7.2 implies a sequence of indeterminate edges of strictly increasing weight from this edge to one adjacent to $\hat{v}$.
Consider two adjacent edges $(\tau_{i},\tau_{j}),(\tau^{\prime}_{i},\tau^{\prime}_{j})$ in this chain, send at times $t_{i},t_{j},t_{i}^{\prime},t_{j}^{\prime}$, respectively. By Lemma 7.2, it must be the case that $t_{j}-\Delta\leq t_{j}^{\prime}$. Adding this bound over each link in the chain shows that $T-2\Delta-(k-1)\Delta-\Delta\leq t$ (where the last $\Delta$ comes from bounding the delay of the transaction in the last edge adjacent to $\hat{v}$).
A sequence can be of length at most $n-1$.
As such, after time $(n+1)\Delta$, every edge adjacent to $\tau$ is either rejected or included and marked determinate.
To ensure that $\tau$ is included in the output of Algorithm 3, it suffices to ensure that every transaction adjacent to an indeterminate edge must come after $\tau$. This is guaranteed by waiting for an additional $\Delta$ time. ∎
7.3. Trading Accuracy for Liveness
The $O(n\Delta)$ bound in Theorem 7.2 comes from the fact that a chain of indeterminate edges can have $O(n)$ length, which, in turn, follows from the fact that weights have a granularity of $\frac{1}{n}$. Reducing this granularity by rounding weights reduces the maximum length of a chain of indeterminate edges.
{lemmarep}
If edge weights are rounded to the nearest $\frac{1}{k}$ in the streamed ordering graph (before Algorithm 3 is applied) then a transaction $\tau$ is contained in the output after at most $(k+2)\Delta$ time.
Proof.
The argument proceeds exactly as in that of Theorem 7.2, except that the length of the chain of indeterminate edges is at most $k$. ∎
However, rounding weakens fairness guarantees.
{lemmarep}
Rounding edge weights to the nearest $\frac{1}{k}$ before applying Algorithm 3 achieves $(\gamma,\frac{f}{n}+\frac{1}{2k})$-minimal-batch-order-fairness for all $\gamma$ and $f$ faulty replicas.
Proof.
The argument proceeds as in Theorem 5. However, due to the rounding and the faulty replicas, the weight on an edge observed by the algorithm might be up to $\frac{f}{n}+\frac{1}{2k}$ different from the true observed weight (as opposed to $\frac{f}{n}$, as in Theorem 5). ∎
8. Protocol Instantiation
The algorithms thus far have assumed that every replica will eventually vote on every transaction. Any faulty replica, however, could exclude a transaction from its output, leaving the transaction (and the protocol) permanently stuck.
The main idea to resolve this difficulty is that if a replica refuses to vote on a transaction, then that replica must be faulty. If a replica is faulty, then it suffices for the rest of the replicas to make up a vote on its behalf.
We give here an example method for instantiating our sequencing protocol on top of existing protocol components, although this is far from the only method.
As a base layer, we require some consensus protocol, such as (yin:hotstuff; lokhava:stellar; castro1999practical; miller2016honey), that proceeds in rounds. At any time, every replica can submit a continuation of its previous ordering; that is, an ordered list of transactions that extends its ordering vote. The consensus protocol then comes to consensus on a set of received ordering votes. The Streaming Ranked Vote algorithm is run (by every replica) on the set of received votes after every new set of votes is finalized by the consensus protocol. Assuming correctness of the consensus protocol, every replica will run the algorithm on the same input, and thereby produce the same output ordering.
We assume that every replica has a registered public key. Every vote continuation must be cryptographically signed by the associated secret key. Furthermore, every continuation must be accompanied by the hash of the previously submitted continuation, so that if one continuation vote from a replica is lost (perhaps by malicious activity of consensus participants) the replica’s ordering cannot be manipulated.
We make a network synchrony assumption, akin to Assumption 7.2. If a transaction appears first in some replica’s vote in round $T$ of the consensus protocol, then we assume that there is some known bound $k$ such that every other honest replica will have time to receive and submit a vote on the transaction before the end of round $T+k$. As such, after every round $r$, the protocol computes the set of all transactions that have not received votes from every replica but were voted on by some replica at or before round $r-k$. For each of these transactions $\tau$, the protocol appends a vote for $\tau$ at the end of a replica’s current vote sequence if $\tau$ is not already present. When appending multiple transactions to a replica’s vote, transactions can be sorted deterministically (i.e. by hash).
This protocol guarantees that every transaction will receive a vote from all replicas within at most $k$ rounds after it is first submitted to the protocol. In this setting, $k$ must be configured in advance by the protocol. The parameter must be set so that every honest replica is able to submit a vote on a transaction within $k$ rounds. Naturally, the choice of $k$ depends on network conditions the censorship-resistance (or “chain-quality” (garay2015bitcoin)) of the consensus protocol.
As an example alternative, one could also use the protocol in kelkar2020order, which is based on atomic broadcast and a “Set Byzantine Agreement” primitive. This protocol waits until all honest replicas are able to add a transaction $\tau$ to their ordering votes, and then invokes a protocol to agree on the set of replicas $U_{\tau}$ which have the transaction in their ordering votes. This set may not be the set of all replicas, but includes all honest replicas. To use our ordering protocol on this consensus protocol, it suffices to make up a vote for the (presumed faulty) replicas not in $U_{\tau}\cap U_{\tau^{\prime}}$. Replicas in $U_{\tau}\setminus U_{\tau^{\prime}}$ can be presumed to vote $\tau\prec\tau^{\prime}$ and vice versa, and replicas not in $U_{\tau}\cup U_{\tau^{\prime}}$ can be presumed to vote (arbitrarily) by, e.g., comparing deterministic hashes of $\tau$ and $\tau^{\prime}$.
9. Conclusion
We study here the problem of ordering transactions in a replicated state machine as a novel streaming instance of the preference aggregation problem from classical social choice theory. This viewpoint enables us to strengthen the notions of “order-fairness” used in prior work, and then construct algorithms solving this problem with both strictly stronger “fairness” guarantees and strictly stronger liveness properties than all of the prior work. To be specific, our streaming variant of Ranked Pairs satisfies $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness for every $\frac{1}{2}<\gamma\leq 1$ simultaneously, and for any number of faulty replicas $f$. Fairness guarantees smoothly weaken as the number of faulty replicas increases. For comparison, prior work must fix a choice of $\gamma$ and a bound on the number of faulty replicas in advance, and can only satisfy $\gamma$-batch-order-fairness for that $\gamma$.
These notions of “order-fairness” are not the only desiderata with practical significance. Different contexts in which a system is deployed pose different constraints and financial incentives. We believe that social-choice style methods for preference aggregation raises many interesting open questions of practical import. For example, what aggregation rules lead to high social welfare? If clients bribe replicas to prefer certain orderings, what rules maximize (or minimize) this fee revenue? How do incentives distort the output ordering?

### v4

@@
-Fair Ordering via Streaming Social Choice Theory
+Fair Ordering in Replicated Systems via Streaming Social Choice
 Abstract.
-Prior work (themis; kelkar2020order; cachin2022quick; vafadar2023condorcet) studies the question of “fairly” ordering transactions in a replicated state machine.
-Each of $n$ replicas receives transactions over a network in a possibly different order, and the system must aggregate these observed orderings into a single order.
-We argue that this problem is best viewed through the lens of social choice theory, in which (in the preference aggregation problem) rankings on candidates are aggregated into an election result.
+How can we order transactions “fairly” in a replicated state machine (of which today’s blockchains are a prototypical example)?
+In the model of prior work (themis; kelkar2020order; cachin2022quick; vafadar2023condorcet; kiayias2024ordering), each of $n$ replicas observes transactions in a different order, and the system aggregates these observed orderings into a single order.
+We argue that this problem is best viewed directly through the lens of the classic preference aggregation problem of social choice theory (instead of as a distributed computing problem), in which rankings on candidates are aggregated into an election result.
 Two features make this problem novel and distinct.
@@
 And second, decisions must be made quickly and with only partial information.
-Additionally, some faulty replicas might alter their reported observations; their influence on the output should be bounded and well understood.
-The prior work studies a “$\gamma$-batch-order-fairness” property, which divides an ordering into contiguous batches.
+Additionally, some faulty replicas might misreport their observations; the influence of faulty replicas on the output should be well understood.
+Prior work studies a “$\gamma$-batch-order-fairness” property, which divides an ordering into contiguous batches.
 If a $\gamma$ fraction of replicas receive a transaction $\tau$ before another transaction $\tau^{\prime}$, then $\tau^{\prime}$ cannot be in an earlier batch than $\tau$.
-We strengthen this definition to require that batches have minimal size ($\gamma$-batch-order-fairness can be vacuously satisfied by arbitrarily large batches) while accounting for the possibility of faulty replicas.
-The lens of social choice theory allows us to give an ordering protocol with both strictly stronger fairness guarantees and better liveness properties than prior work.
-We study the Ranked Pairs method.
-Analysis of how missing information moves through the algorithm allows our streaming version to know when it can output a transaction.
-Deliberate construction of a tiebreaking rule ensures our algorithm outputs a transaction after a bounded time (in a synchronous network).
-Prior work relies on a fixed choice of $\gamma$ and bound on the number of faulty replicas $f$, but we show that Ranked Pairs satisfies our definition for every $\frac{1}{2}<\gamma\leq 1$ simultaneously and for any $f$, where fairness guarantees linearly degrade as $f$ increases.
+This definition holds vacuously, so we strengthen it to require that batches have minimal size, while accounting for faulty replicas.
+This lens gives a protocol with both strictly stronger fairness and better liveness properties than prior work.
+We specifically adapt the Ranked Pairs (tideman1987independence) method to this streaming setting.
+This algorithm can be applied on top of any of the communication protocols (in various network models) of prior work for immediate liveness and fairness improvements.
+Prior work relies on a fixed choice of $\gamma$ and a bound on the number of faulty replicas $f$, but we show that Ranked Pairs satisfies our definition for every $\frac{1}{2}<\gamma\leq 1$ simultaneously and for any $f$, where fairness guarantees degrade as $f$ increases.
 copyright: none
@@
 An (unbounded) set of clients create and broadcast new transactions.
-A classic example is a banking application, in which the state machine maintains a list of user account balances, and a transaction transfers money from one user to another.
+This architecture famously underlies many of today’s blockchains, but also underlies many traditional (centralized) services, such as bank infrastructure.
 This architecture requires that all replicas apply transactions in the same order.
-In the banking example, a client $A$ with a \$1 balance might create two transactions, one which sends a \$1 payment to client $B$ and one which sends a \$1 payment to client $C$.
+In a bank example, a client $A$ with a \$1 balance might create two transactions, one which sends a \$1 payment to client $B$ and one which sends a \$1 payment to client $C$.
 Only one of payments can execute successfully (assuming $A$’s balance is not allowed to overdraft) so if different replicas apply these transactions in different orders, then they will disagree on the balances of clients $B$ and $C$.
@@
 The key observation for this work is that this problem is a novel streaming variation of the classic preference aggregation problem of social choice theory (hagele2001lulls; llull; colomer2013ramon; condorcet1785essay; arrow1950difficulty).
-Prior work on this problem (kelkar2020order; themis; cachin2022quick) focuses on network communication, but (through varying protocols) each produces agreement on a “vote” from each replica on the ordering of a set of transactions (in social choice parlance, a “ranking” on a set of “candidates”).
+Prior work on this problem (kelkar2020order; themis; cachin2022quick; aequitaspermissionless) use very different network communication protocols, but ultimately each produces agreement on a “vote” from each replica on the ordering of a set of transactions (in social choice parlance, a “ranking” on a set of “candidates”).
 Agnostic to the choice of network protocol, how should a system aggregate a set of proposed orders (“rankings”) into a single, total order?
@@
 This algorithm could be deployed on top of any of the network protocols of prior work.
-As an additional complication, replicas might adjust their reported orderings in response to strategic considerations.
-One desirable property of an order aggregation rule is that the influence of a (potentially colluding) subset of replicas is bounded and precisely quantified.
-Mirroring the prior work, we write here that it is “faulty” behavior for a replica to report anything other than the order in which it observes transactions arriving over the network.
+One additional consideration is that replicas might strategically adjust their reported orderings.
+Prior work (kelkar2020order; themis; cachin2022quick) writes that “honest” replicas must report transactions in the order in which they arrive over a network, and any other behavior is “faulty.” One desirable property of an order aggregation rule is that the influence of a (colluding) subset of faulty replicas is precisely bounded.
 There is a wide body of social choice literature on this underlying aggregation problem.
@@
 1.1. Our Results
-To the best of our knowledge, this work is the first to directly pose the question as a streaming version of the classic preference aggregation problem.
 We first observe that for an ordering method to be well-defined on a countably infinite set, it suffices for there to exist a “monotonic” and “asymptotically live” algorithm that implements the ordering method on finite, initial segments of the input.
@@
 [^1] Our algorithms do not require knowledge of a bound on the number of faulty replicas.
-This number may be separate from a bound on the number of faulty replicas in an underlying communication protocol.
-Standard Byzantine-fault tolerant communication protocols require $n\geq 3f+1$.
+This notion of “faulty” replica is distinct from notions of “faulty” in classical Byzantine-fault tolerant communication protocols, such as (castro1999practical).
 This work adapts the ordering known as Ranked Pairs (tideman1987independence) to our streaming setting.
@@
 Ranked Pairs iterates over edges in order of edge weight.
-We find that the tiebreaking rule in this ordering is a critical part of the algorithm.
-Using a naive deterministic rule, our streaming algorithm might never output any transactions.
-However, we then show in §7 that deliberate construction of the tiebreaking rule guarantees asymptotic liveness.
-Finally, a synchronous network assumption combined with our tiebreaking guarantees that our algorithm outputs each transaction after a finite amount of time.
-If at most $\Delta$ time elapses between when a transaction $\tau$ is sent and when every honest replica votes on it, for example, our algorithm outputs $\tau$ after a delay of at most $O(n\Delta)$.
+An interesting observation is that the tiebreaking rule in this edge ordering critically determines overall liveness.
+A worst-case (fixed) rule could cause our streaming algorithm to never output any transactions.
+However, we show in §6 how to construct a tiebreaking rule in a streaming fashion that guarantees not only asymptotic liveness but also an explicitly bounded liveness delay.
+If, for example, at most $\Delta$ time elapses between when a transaction $\tau$ is sent and when every honest replica votes on it, for example, our algorithm outputs $\tau$ after a delay of at most $O(n\Delta)$.
 This can be reduced to $O(k\Delta)$, by rounding edge weights to the nearest $\frac{1}{k}$— although this rounding reduces the fairness guarantee to $(\gamma,\frac{f}{n}+\frac{1}{2k})$-minimal-batch-order-fairness (again for every $\frac{1}{2}<\gamma\leq 1$ simultaneously, and any $f$).
@@
 Otherwise, the replica is faulty.
-We assume that at most $f$ of the $n$ replicas are faulty, although our algorithms do not require knowledge of $f$.
+We denote the number of faulty replicas as $f$ (out of $n$), and do not assume a bound or knowledge of $f$.
 333 Of course, $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness (Definition 3.4) is only meaningful if $f<\frac{n}{2}$.
@@
 A precise choice of response to this type of faulty behavior will depend on the network model and communication protocol, but, informally, if a replica fails to vote on a transaction within a “sufficient” time, it suffices for the honest replicas to fabricate a vote on behalf of the faulty replica.
-We address this difficulty in §8.
-3. Related Work
-3.1. Order-Fairness
-Closest to this work are Aequitas (kelkar2020order) and Themis (themis), which propose the definition of $\gamma$-batch-order-fairness and instantiate ordering protocols based on aggregating ordering votes into a total order.
-Aequitas achieves $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness for a fixed choice of $\gamma$, but is not asymptotically live.
-§4.3 gives a more detailed comparison to these works.
-Cachin et al.
-(cachin2022quick) gives a notion of “$\kappa$-differential-order-fairness;” Theorem D.1 of (themis) shows an equivalence between this definition and $\gamma$-batch-order-fairness, for a particular choice of $\gamma$ (as a function of $\kappa$).
-Our definition of $(\gamma,\delta)$-minimal-batch-order-fairness (Definition 4.4) strengthens the (vacuously satisfiable) $\gamma$-batch-order-fairness definition used in prior work.
-§4 discusses the details of this relationship.
-Concurrently with this work, vafadar2023condorcet observe that an adversary can influence the output ordering of many protocols (including that of kelkar2020order) by sending extra transactions (which can force batches to be combined together).
-This eliminates the constraints on the output that Definition 4.1 gives (using an example akin to Example 11.2).
-Lemma 4.3 identifies a similar problem.
-This observation is a corollary of the fact that the protocol of kelkar2020order (as well as the Ranked Pairs algorithm studied here) does not satisfy the Irrelevance of Independent Alternatives axiom (arrow1950difficulty).
-vafadar2023condorcet also propose using Ranked Pairs to sort transactions within batches (as, e.g., those output by a protocol like Themis (themis)).
-However, no modification to the ordering of transactions within a batch output in Aequitas (kelkar2020order) can give the Ranked Pairs ordering, because Ranked Pairs might need to interleave batches that Aequitas considers incomparable (as demonstrated in §12).
-This strategy also cannot mitigate Aequitas’s lack of asymptotic liveness.
-This mitigation, if applied to Aequitas, is akin to dropping all edges of weight less than $\gamma$ before running Ranked Pairs.
-This strategy, therefore, would achieve $(\gamma^{\prime},\frac{f}{n})$-minimal-batch-order-fairness for all $\gamma^{\prime}\geq\gamma$ (a weaker guarantee than what our protocol provides).
-As discussed in §4.3, the batches in Themis (themis) are defined slightly differently than in Definition 4.1.
-Themis, when computing a graph of ordering dependencies, adds a directed edge between every pair of vertices, choosing the direction of an edge by majority vote (ignoring the parameter $\gamma$, which is considered only on edges adjacent to transactions that some replicas have not yet reported observing).
-In so doing, it merges together any two batches unless, for any transaction $\tau$ in one and $\tau^{\prime}$ in the other, a majority of replicas vote $\tau\prec\tau^{\prime}$ (which bypasses the counterexample of §12).
-Additionally, Themis’s (themis) round-based sequencing protocol presents a different problem insurmountable by any process like the one suggested by vafadar2023condorcet that sorts batches separately.
-Namely, Themis may finalize the composition of a batch before has enough information to do guarantee that, according to the Ranked Pairs ordering, all transactions in the batch must come before all those not in the batch (and not already in the output).
-This means that Themis may output two batches in sequence, such that running Ranked Pairs on each batch separately produces a different output than running Ranked Pairs on the union of the two batches together.
-We give a concrete example of this phenomenon in §13.
-aequitaspermissionless instantiate Aequitas in a setting where the number of replicas is not known.
-zhang2020byzantine adds timestamps to ordering votes and sorts transactions by median reported timestamp to provide a linearizability guarantee between transactions.
-This guarantee is incomparable to $(\gamma,\delta)$-minimal-batch-order-fairness.
-3.2. Social Choice
-Classical work of social choice studies the problem of choosing an order between a fixed, finite set of candidates given a complete set of preferences from each voter (arrow1950difficulty).
-Prior work studies related models in a limited-information setting.
-Lu and Boutilier (lu2013multi), Ackerman et.
-al (ackerman2013elections), and Cullinan et.
-al (cullinan2014borda) study the setting where a social choice rule has incomplete access (a partial ordering) to a voter’s preferences over a finite set of candidates.
-Conitzer and Sandholm (conitzer2005communication) study the communication complexity of a variety of voting rules.
-Fain et.
-al (fain2019random) study a randomized ordering rule that requires only a constant number of queries to voter preferences.
-Fishburn (fishburn1970arrow) showed that Arrow’s impossibility theorem does not hold when the set of voters is infinite, although Kirman and Sondermann (kirman1972arrow) find “dictatorial sets” of voters.
-Grafe and Grafe (grafe1983arrow) extend these results to the case of infinitely many alternatives, subject to a continuity condition on the space of ranking preferences.
-Chichilnisky and Heal (chichilnisky1997social) and Efimov and Koshevoy (efimov1994topological) study the types of rules admissible in the infinite voters setting, given a topology on ordering preferences.
-crisman2022voting study the preference aggregation problem where the output must be a cyclic permutation of the candidates.
-3.3. Front-Running
-The study of order fairness was initially motivated (as in, e.g., kelkar2020order) by the goal of limiting an adversary’s ability to order transactions (“front-running”) in public blockchains for their own financial benefit (daian2020flash).
-Additional approaches to this front-running problem include commiting to a transaction ordering before any replica can know transaction contents, using threshold encryption or commit-reveal schemes (malkhi2022maximal; clineclockwork; zhang2022flash).
-Li et.
-al (li2023transaction) study the problem of verifiably computing an ordering within an encrypted hardware enclave (costan2016intel).
-These could be applied on top of a streaming ordering algorithm.
-Kavousi et.
-al (kavousi2023blindperm) randomly shuffle blocks of encrypted transactions.
-Some decentralized exchanges (ramseyer2023speedex; cowswapproblem; penumbraswap) process transactions in unordered batches, thereby eliminating the need to compute an ordering.
-ferreira2022credible construct an exchange-specific ordering that provides traders with certain worst-case price guarantees.
-mamageishvili2023buying use a median-timestamp mechanism akin to zhang2020byzantine, but allow clients to pay to reduce its reported timestamp.
+§16 gives an example construction that satisfies this property, but note that prior work on this problem (e.g., (kelkar2020order; themis; cachin2022quick)) implicitly use the same ideas to handle unresponsive replicas.
 3. Defining Fair Ordering
 3.1. Batch Order Fairness
-Our discussion here follows that of (kelkar2020order; themis; cachin2022quick), but gives, unlike prior work, a definition that is meaningful and achievable.
-Honest replicas might receive transactions over the network in a different order.
-To resolve disagreements “fairly,” recent work (kelkar2020order) proposes a notion known as “$\gamma$-Batch-Order-Fairness,” which we reproduce below from Definition III.1 of themis (a more complicated but functionally similar definition is used in kelkar2020order).
+Our discussion here follows the notion of “fairness” discussed in prior work (e.g., (kelkar2020order; themis; cachin2022quick; kiayias2024ordering)).
+The natural notion of “fairness” in prior work, which we follow here, is that if “most” replicas report a transaction $\tau$ before another transaction $\tau^{\prime}$, then the protocol outputs $\tau$ before $\tau^{\prime}$.
+Making this notion precise requires accounting for what are known as Condorcet cycles.
+A (super)majority might report $\tau$ before $\tau^{\prime}$, another majority reports $\tau^{\prime}$ before $\tau^{\prime\prime}$, and a third majority reports $\tau^{\prime\prime}$ before $\tau$.
+Aequitas (kelkar2020order) sidesteps this problem by outputting these Condorcet cycles in “batches,” which leads to the following definition (paraphrased from (kelkar2020order)).
 Definition 3.1 ($\gamma$-batch-order-fairness).
-Suppose that $\tau$ and $\tau^{\prime}$ are received by all nodes. If $\gamma n$ nodes received $\tau$ before $\tau^{\prime}$, then a ranking algorithm outputs $\tau$ no later than $\tau^{\prime}$.
-Aequitas (kelkar2020order) outputs transactions in batches, and “no later than” means “in the same batch.”Transactions within a batch are ordered arbitrarily.
-Themis (themis) outputs a total ordering, with the assertion that this total ordering could be divided into contiguous (disjoint) batches satisfying this property.
-One challenge is the parameter $\gamma$.
-It is not obvious if a higher or lower $\gamma$ gives stronger fairness properties.
-Indeed, there are simple examples (with no faulty replicas) where Definition 4.1 (actually the stronger Definition 4.4) only implies ordering restrictions at high values of $\gamma$, others at only low values, and still others at only intermediate values (§11).
-Prior work relies on a fixed choice of $\gamma$, but our algorithms satisfy Definition 4.4 for all values of $\gamma$ simultaneously.
+Suppose that $\tau$ and $\tau^{\prime}$ are received by all nodes.
+If $\gamma n$ nodes received $\tau$ before $\tau^{\prime}$, then a ranking algorithm never outputs $\tau^{\prime}$ in a later batch than $\tau$.
+In other words, $\tau$ and $\tau^{\prime}$ could be in the same batch.
+Transactions within a batch are ordered arbitrarily.
+Subsequent work (themis; cachin2022quick; kiayias2024ordering) follows this same definition (or a restatement of it).
+There are two key problems with this definition.
+First, this definition does not preclude a ranking algorithm from putting all transactions into one batch (and therefore ordering transactions arbitrarily).
+For this definition to be meaningful, it needs a notion of “minimality” to batch sizes.
+We address this problem by strengthening the definition to include a notion of “minimality” (Definition 3.4).
+Some of the prior work (kelkar2020order; cachin2022quick; kiayias2024ordering), but not all (themis), implicitly satisfies this definition.
+Second, it is not clear whether a higher or lower $\gamma$ gives stronger fairness properties, yet all of the prior work requires a fixed parameter choice.
+Indeed, there are simple examples (with no faulty replicas) where Definition 3.4 only implies ordering restrictions at high values of $\gamma$, others at only low values, and still others at only intermediate values (§10).
+It is not the case that satisfying Definition 3.4 for a low value of $\gamma$ implies satisfying the definition for a high value of $\gamma$.
+We address this problem by constructing a protocol that satisfies Definition 3.4 for all $\frac{1}{2}\leq\gamma\leq 1$ simultaneously.
 3.2. Batch Minimality
-Unfortunately, Definition 4.1 is insufficient to provide meaningful ordering guarantees.
-A protocol that outputs all transactions in the same batch vacuously satisfies this definition, and if transactions within a batch can be ordered arbitrarily, then every ordering satisfies Definition 4.1 for every $\gamma$.
-Prior work (themis) recognizes the need for “minimality”.However, the notion of minimality in prior work is not achievable in the presence of faulty replicas (Theorem 4.2).
-Towards a feasible notion of minimality, we first define a graph denoting ordering dependencies.
+To construct a precise definition, we start with a weighted dependency graph over a set of transactions.
 Definition 3.2 (Ordering Graph).
 Suppose that replicas report ordering votes $(\sigma_{1},...,\sigma_{n})$ on a set of transactions $V$.
-The Ordering Graph $G(\sigma_{1},...\sigma_{n})$ is a complete, weighted directed graph with vertex set $V$ and, for each transaction pair $(\tau,\tau^{\prime})$, an edge of weight $w(\tau,\tau^{\prime})=\alpha$ if $\alpha n$ replicas report receiving $\tau$ before $\tau^{\prime}$.
+The Ordering Graph $G(\sigma_{1},...\sigma_{n})$ is a complete, weighted directed graph with vertex set $V$ and, for each transaction pair $(\tau,\tau^{\prime})$, an edge of weight $w(\tau,\tau^{\prime})=\alpha$ if $\alpha n$ replicas report $\tau\prec\tau^{\prime}$.
 Any notion of “minimal” should at least require that when there are no cycles in ordering dependencies, the output is consistent with every ordering dependency (as in Theorem IV.1 part 1, Themis (themis)).
@@
 For any pair of transactions $\tau,\tau^{\prime}$, if $\gamma n$ replicas receive $\tau\prec\tau^{\prime}$ but $\tau^{\prime}$ is output before $\tau$, then there exists a sequence of transactions $\tau^{\prime}=\tau_{1},...,\tau_{k}=\tau$ where $\gamma n$ replicas receive $\tau_{i}\prec\tau_{i+1}$ and $\tau_{i}$ is output before $\tau_{i+1}$ for all $i$.
-Unfortunately, this notion of minimality is impossible to achieve (for any $\gamma$) in the face of any faulty replicas.
-The condition on $f$ in Theorem 4.2 is a technical artifact of a construction in the proof (related to the pigeonhole principle).
+However, this notion of minimality is impossible to achieve (for any $\gamma$) in the face of any faulty replicas.
+The condition on $f$ in Lemma 3.2 is a technical artifact of a construction in the proof (related to the pigeonhole principle).
 For some parameter settings, such as $n=7$ with $\gamma=\frac{2}{3}$, the condition admits $f=1$.
-{theoremrep}
+{lemmarep}
 No protocol can achieve $\gamma$-batch-order-fairness with exactly $\gamma$-minimal output batches for any $n$ and $f$ greater than $n\mod(n-\lceil\gamma n\rceil+1)$ (for $\gamma>1/2$).
@@
 ∎
-There is one subtle difference between this definition and (an analogue of) the notion of “minimality” in Theorem IV.1 of Themis, and the sequence of batches output in Aequitas.
+There is one subtle difference between this definition and the requirement for an explicit sequence of output batches in prior work (e.g., in (kelkar2020order; themis)).
 When there are two disjoint strongly connected components with no dependencies (of strength at least $\gamma$) from one to the other, Definition 3.4 allows the output to interleave these components.
@@
 Before constructing our protocol, we first apply Definition 3.4 to related prior work.
+Aequitas (kelkar2020order) is instantiated with a choice of parameter $\gamma$ and a bound $f$ on the number of faulty replicas.
 {lemmarep}
-Aequitas with parameters $(\gamma,f)$ achieves $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness.
+Aequitas (kelkar2020order) with parameters $(\gamma,f)$ achieves $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness.
 Note that Aequitas requires $\gamma-\frac{1}{2}>\frac{2f}{n}$. Lemma 3.3 is a stronger version of Theorem 6.6 of (kelkar2020order).
@@
 ∎
+Broadly speaking, Aequitas, along with the protocols of cachin2022quick and kiayias2024ordering, follow the same pattern.
+Through (very different) communication protocols, they come to agreement on, for each transaction pair $\tau$ and $\tau^{\prime}$, the fraction of replicas that report $\tau$ before $\tau^{\prime}$.
+These protocols choose (in advance) a fixed parameter $\gamma$, and then drop from consideration ordering dependences of strength less than $\gamma$.
+This is equivalent to constructing the ordering dependency graph of Definition 3.2, but then dropping all edges of weight less than $\gamma$ and then forgetting all of the weights on the remaining edges.
+The strongly connected components of this graph form the batches output in Aequitas.
+Aequitas (as well as the protocols of cachin2022quick and kiayias2024ordering) also is not asymptotically live (Definition 5.2).
+The strongly connected components of this unweighted dependency graph can be arbitrarily large, and these protocols wait until they observe the entirety of a component before outputting any transaction in it.
+An exception is one version of the protocol in kiayias2024ordering, which adds timestamps to transactions and (given precise network assumptions) can output a component in a streaming fashion.
 By contrast, Ranked Pairs (Theorem 4) achieves $(\gamma,\frac{f}{n})$-minimal-batch-fairness for every $\gamma$ simultaneously, and for any $f$ (our work does not require knowledge of a bound on $f$).
-Aequitas also is not asymptotically live (Definition 6.2).
-Themis claims (Theorem IV.1, (themis)) to output a sequence that can be partitioned into minimal batches, which appears to contradict Theorem 4.2.
-However, Themis relaxes its notion of an “ordering dependency”.
+Our protocol gets these stronger results by using all of the information available in the problem input.
+Additionally, Ranked Pairs guarantees a bounded liveness delay (although precise end-to-end results depend also on network conditions).
+Themis (themis) claims (Theorem IV.1) that its output can be partitioned into “minimal” batches, which appears to contradict Lemma 3.2. However, Themis uses a much weaker notion of “minimal.”
 Where $\gamma$-batch-order-fairness would create a dependency for $\tau$ not after $\tau^{\prime}$ if at least $\gamma n$ replicas receive $\tau$ before $\tau^{\prime}$, the relaxation implied by Definition III.1 of (themis) creates such a dependency if at least $n(1-\gamma)+1$ receive $\tau$ before $\tau^{\prime}$ (that is, it cannot be the case that more than $\gamma n$ receive $\tau^{\prime}$ before $\tau$).
-Under this definition, the construction in Theorem 4.2 constitutes what the authors of (themis) call a “weak Concorcet Cycle”.
-Extra ordering dependencies, however, can artificially merge batches together, giving counterintuitive results.
-{lemmarep} In Themis (themis), even if all nodes report $\tau$ before $\tau^{\prime}$ and there are no faulty replicas nor cycles in ordering dependencies (as defined in Definition 4.1) for any $\gamma>1/2$, the output may put $\tau^{\prime}$ before $\tau$.
-Lemma 4.3 follows from the fact that when two transactions $\tau,\tau^{\prime}$ receive $n-f$ votes in one round of Themis, it determines whether to consider as a dependency $\tau\prec\tau^{\prime}$ or $\tau^{\prime}\prec\tau$ by majority vote (disregarding $\gamma$).
-The minimality statement of Theorem IV.1 of (themis) is with regard to this weaker notion of dependency.
+Theorem IV.1 of (themis) considers its output batches “minimal” if they are singletons when there are no cycles of these dependencies.
+This notion of minimality is much weaker than what we discuss here, and does not provide any guarantee when there is significant disagreement between replicas on transaction orderings.
+Lemma 3.3 gives an example highlighting this distinction.
+{lemmarep}
+In Themis (themis), even if all nodes report $\tau$ before $\tau^{\prime}$ and there are no faulty replicas nor cycles in ordering dependencies (as defined in Definition 3.1) for any $\gamma>1/2$, the output may put $\tau^{\prime}$ before $\tau$.
+This kind of counterintuitive behavior is a possible output of Themis’s protocol.
+When two transactions $\tau,\tau^{\prime}$ receive $n-f$ votes in one round of Themis, it determines whether to consider as an ordering dependency $\tau\prec\tau^{\prime}$ or $\tau^{\prime}\prec\tau$ by majority vote.
+Equivalently, Themis as a protocol operates in the regime of $\gamma=\frac{1}{2}$, regardless of a chosen parameter value (the choice of $\gamma$ does affect how Themis considers transactions reported by some but not all replicas, which does not affect this discussion).
 Proof.
@@
 4. Ranked Pairs Voting
-We turn now to the Ranked Pairs method (tideman1987independence). Ties in edge weights are broken according to an arbitrary (deterministic) rule.
-Input: An ordering vote from each replica $\{\sigma_{i}\}_{i\in[n]}$
-$G=(V,E,w)\leftarrow G(\sigma_{1},...,\sigma_{n})$ ;
-  /* Compute ordering graph */
-$H\leftarrow(V,\emptyset)$ ;
-  /* Initialize empty graph on same vertex set */
-Sort $E$ by edge weight
-foreach $e\in E$ do
-       Add $e$ to $H$ if it would not create a directed cycle in $H$
-end foreach
-Output: The topological sort of $H$
-  /* The loop considered every edge, so the topological sort of $H$ is unique */
-ALGORITHM 1 Ranked Pairs Voting
+Our protocol is based on the Ranked Pairs method (tideman1987independence).
+Ranked Pairs (Algorithm 2, reproduced in §15) first builds the ordering graph (Definition 3.2), then iterates through edges in order of edge weight (breaking ties arbitrarily).
+It builds an acyclic set of edges greedily, adding each new edge so long as it does not create a cycle with edges already in the set.
+The output ordering is the result of topologically sorting the resulting acyclic graph.
 {theoremrep}
@@
 Lemma 3.2 and Theorem 4 together imply an important property.
-The output of Ranked Pairs can be divided into batches consistent with $\gamma$-batch-order-fairness for any $\gamma$, not just a single $\gamma$ (subject to the interleaving discussed in §4.2).
-Unlike prior work, there is no need to choose a $\gamma$ or order transactions within a batch.
+Ranked Pair’s output can be divided into batches consistent with $\gamma$-batch-order-fairness for any $\gamma$ (subject to the interleaving discussed in §3.2).
+There is no need to choose a $\gamma$ or arbitrarily order transactions within a batch.
 Proof.
@@
 In a real-world system, such an algorithm would be run with some periodic frequency, and replicas would periodically append to their ordering votes.
-Our algorithms must be monotonic (Definition 6.1)—if replicas extend their ordering votes, the algorithm, when run again on the extensions, can only extend its prior output.
+An algorithm must be monotonic (Definition 5.1)—if replicas extend their ordering votes, the algorithm, when run again on the extensions, can only extend its prior output.
 Definition 5.1 (Monotonicity).
 A sequencing algorithm $\mathcal{A}(\cdot,\ldots,\cdot)$ is monotonic if, given two sets of ordering votes $\left(\sigma_{1},\ldots,\sigma_{n}\right)$ and $\left(\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{n}\right)$, such that $\sigma^{\prime}_{i}$ extends $\sigma_{i}$ for all $i\in[n]$, $\mathcal{A}(\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{n})$ extends $\mathcal{A}(\sigma_{1},\ldots,\sigma_{n})$.
-Observe that if a sequencing algorithm is monotonic, it implies a well-formed definition for aggregating a set of orderings on countably infinite sets of transactions, not just on finite inputs.
-Specifically, $\tau$ comes before $\tau^{\prime}$ in the overall (infinite) case if there exists a finite subset of the input orderings such that the algorithm puts $\tau\prec\tau^{\prime}$ in its output.
+If a sequencing algorithm is monotonic, it implies a well-formed definition for aggregating a set of orderings on countably infinite sets of transactions, not just on finite sets.
+$\tau$ comes before $\tau^{\prime}$ in the infinite case if there exists a finite subset of the input orderings such that the algorithm puts $\tau\prec\tau^{\prime}$ in its output.
 A non-vacuous condition is also required—the algorithm that never outputs anything is monotonic.
@@
 Furthermore, the set of vertices that such a path can visit are bounded to a local neighborhood of $\tau_{A}$ and $\tau_{B}$.
-Definition 6.3 ().
+Definition 5.3.
 Let $\tau$ be any transaction.
@@
 Proof.
-If the path visits a transaction $\tau$ in $R_{\tau_{A}}$, then it creates a cycle with the edge from $\tau_{A}$ to $\tau$ (which ranked pairs must choose, by Observation 6.1).
+If the path visits a transaction $\tau$ in $R_{\tau_{A}}$, then it creates a cycle with the edge from $\tau_{A}$ to $\tau$ (which ranked pairs must choose, by Observation 5.1).
 If the path visits a transaction $\tau$ in $P_{\tau_{B}}$, then it creates a cycle with the edge from $\tau$ to $\tau_{B}$.
@@
 As such, we allow our algorithms to leave an edge in an “indeterminate” state, and design our algorithms to account for indeterminate edges when considering subsequent edges.
-First, we construct an ordering graph that captures the information available to an algorithm at a finite time.
+First, we construct an ordering graph that captures the information available at a finite time.
 Definition 5.4 (Streamed Ordering Graph).
-Suppose that each replica submits an ordering vote $(\sigma_{1},...,\sigma_{n})$. Let $V$ be the set of all transactions that appear in each vote, and let $\hat{v}$ be the “future” vertex.
-The Streamed Ordering Graph is a weighted, directed, complete graph $\hat{G}=(V^{\prime},E)$ on vertex set $V\cup\{\hat{v}\}$.
+Suppose that each replica submits an ordering vote $(\sigma_{1},...,\sigma_{n})$.
+Let $V$ be the set of all transactions that appear in each vote, and let $\hat{v}$ be a new vertex (the “future”) The Streamed Ordering Graph is a weighted, directed, complete graph $\hat{G}=(V^{\prime},E)$ on vertex set $V\cup\{\hat{v}\}$.
 For each $\tau$ and $\tau^{\prime}$, set weights $w(\tau,\tau^{\prime})$ and $w(\tau^{\prime},\tau)$ as in Definition 3.2.
@@
 5.2. A Streaming Algorithm
-We can now construct an algorithm to compute Ranked Pairs in our streaming setting.
-Instead of either including or rejecting every edge, Algorithm 2 has the option to declare an edge as indeterminate; that is, the algorithm is not able to determine whether to include or exclude the edge with the information in its input.
-Input: An ordering vote from each replica $\{\sigma_{i}\}_{i\in[n]}$
-$\hat{G}=(V,E,w)\leftarrow\hat{G}(\sigma_{1},...,\sigma_{n})$ ;
-  /* Compute streamed ordering graph */
-$H\leftarrow(V,\emptyset)$ ;
-  /* Initialize empty graph on same vertex set */
-foreach edge $(\tau,\hat{v})$ and $(\hat{v},\tau)$ with $w(e)=1$ do
-       Add $e$ to $H$, and mark it as indeterminate
-end foreach
-Sort $E$ by edge weight, with a fixed tiebreaking rule.
-foreach $e=(\tau_{i},\tau_{j})\in E$ do
-       $U_{\tau_{i},\tau_{j}}\leftarrow(V\setminus(R_{\tau_{i}}\cup P_{\tau_{j}}))\cup\{\hat{v}\}$
-        /* Recall that $\hat{v}$ is the ‘‘future’’ vertex of the streamed ordering graph */
-       If there is a directed path in $H$ restricted to $U_{\tau_{i},\tau_{j}}$ from $\tau_{j}$ to $\tau_{i}$ where every edge is determinate, then do not include the edge in $H$.
-       If there is no directed path in $H$ restricted to $U_{\tau_{i},\tau_{j}}$ from $\tau_{j}$ to $\tau_{i}$ where every edge is either determinate or indeterminate, add $(\tau_{i},\tau_{j})$ to $H$ and mark it as determinate.
-       Otherwise, add $(\tau_{i},\tau_{j})$ to $H$ and mark it as indeterminate.
-end foreach
-Output: The topological sort of $H$ (with sorting comparisons restricted to determinate edges), up to but not including the first transaction $\tau$ bordering an indeterminate edge
-ALGORITHM 2 Streaming Ranked Pairs
+We can now construct a streaming algorithm to compute Ranked Pairs.
+Instead of either including or rejecting every edge, Algorithm 3 (shown in full in Appendix 15) has the option to declare an edge as indeterminate; that is, the algorithm cannot determine whether to include or exclude the edge with the information in the input.
+The algorithm proceeds identically to the classic Ranked Pairs algorithm, except regarding how it treats indeterminate edges.
+For each new edge, it first checks whether it can accept the edge (without creating a cycle of accepted edges), supposing that all currently indeterminate edges are later accepted.
+Conversely, it also checks whether it can reject the edge, which means checking whether accepting the edge would always create a cycle of accepted edges (assuming the currently indeterminate edges are later rejected).
+The algorithm operates on the modified input graph (Definition 5.4), where the “future” vertex bounds the information the algorithm has to consider about as-yet-unseen transactions.
+Informally, all edges between as-yet-unseen transactions are indeterminate, so from the perspective of directed connectivity analysis, the unseen transactions form a single, strongly-connected component.
 To prove correctness of this algorithm, it suffices to show that the output of the algorithm is monotonic and that the output is consistent with the non-streaming version of Ranked Pairs.
-By consistency, we specifically mean that at any point, if clients were to stop sending new transactions and all replicas eventually voted on every replica, Algorithms 1 and 2 would produce the same output.
+By consistency, we specifically mean that at any point, if clients were to stop sending new transactions and all replicas eventually voted on every replica, Algorithms 2 and 3 would produce the same output.
 {lemmarep}
@@
 Output: The topological sort of $H$ (with sorting comparisons restricted to determinate edges), up to but not including the first transaction bordering an indeterminate edge
-ALGORITHM 3 Streaming Ranked Pairs, Version 2
+ALGORITHM 1 Streaming Ranked Pairs
 6.1. Efficient Marginal Runtime
@@
 6.2. Protocol Liveness
-Algorithm 2 is not asymptotically live. §14 gives an instance on which the algorithm would never output any transaction.
-That construction uses an adversarial tiebreaking rule, but Ranked Pairs requires only some tiebreaking rule.
-Our algorithm can therefore generate a tiebreaking rule dynamically.
+Algorithm 3 is not asymptotically live.
+Appendix 13 gives a worst-case example.
+The problem is the tiebreaking rule for how Ranked Pairs iterates over edges of equal weight.
+That construction uses a worst-case tiebreaking rule, but Ranked Pairs requires only some tiebreaking rule.
+Our algorithm can generate a tiebreaking rule dynamically.
 Specifically, when the algorithm visits edges of weight $\gamma=\frac{k}{n}$, it defers edges that would become indeterminate to the end of the ordering among edges of that weight.
-Algorithm 3 constructs a (partial) ordering between edges that breaks ties between edges of the same weight.
+Algorithm 1 thus constructs a (partial) ordering between edges that breaks ties between edges of the same weight.
 The output of this algorithm is, by the same arguments as in §5.2, consistent with the output of non-streaming Ranked Pairs, when using an tiebreaking rule consistent with this edge ordering.
@@
 ∎
-Crucially, this small change enables the following structural lemma. We start with a useful definition.
-Definition 7.1 ().
+This small change enables the following structural lemma. We start with a useful definition.
+Definition 6.1.
 An edge $(\hat{\tau},\hat{\tau^{\prime}})$ is contemporary to an edge $(\tau,\tau^{\prime})$ if and only if $\hat{\tau}$ and $\hat{\tau^{\prime}}$ are contained in $U_{\tau,\tau^{\prime}}$.
@@
 If Algorithm 1 marks an edge $(\tau_{i},\tau_{j})$ with weight $\frac{k}{n}$ as indeterminate, there must be a path from $\tau_{j}$ to $\tau_{i}$ contained in $U_{\tau_{i},\tau_{j}}$ that contains an indeterminate edge of weight at least $\frac{k+1}{n}$.
-By contrast, the best bound in Algorithm 2 is that there exists a contemporary indeterminate edge of weight $\frac{k}{n}$.
 Proof.
@@
 ∎
+By contrast, the best bound in Algorithm 3 is that there exists a contemporary indeterminate edge of weight $\frac{k}{n}$.
 Lemma 6.2 implies that for any indeterminate edge, there is sequence of contemporary indeterminate edges of strictly increasing weights that causes the edge to be indeterminate.
@@
 ∎
-Additionally, under network synchrony, Lemma 7.2 bounds the temporal delay between when a transaction is sent by a client and when Algorithm 3 adds it to its output.
-Assumption 7.2 (Synchronous Network).
+Additionally, given a network assumption like Assumption 6.2, Lemma 6.2 bounds the temporal delay between when a transaction is sent by a client and when Algorithm 1 adds it to its output.
+Assumption 6.2 (Network Delay Bound).
 If a client sends a transaction $\tau$ at time $t$, all replicas include $\tau$ in their ordering votes before time $t+\Delta$.
-As such, replicas can disagree on the ordering between $\tau$ and $\tau^{\prime}$ only if they were sent at roughly the same time.
+As such, replicas disagree on the ordering between $\tau$ and $\tau^{\prime}$ only if they were sent at roughly the same time.
+Note that Algorithm 1 does not require knowledge of $\Delta$, and does not depend on Assumption 6.2 for correctness.
 {observationrep}
@@
 Proof.
-By construction, $\tau\notin R_{\tau_{i}}$, so $t\leq t_{i}+\Delta$. Furthermore, $\tau\notin P_{\tau_{j}}$, so $t+\Delta\geq t_{j}$. Furthermore, by Observation 7.2, $t_{j}+\Delta\geq t_{i}$.
+By construction, $\tau\notin R_{\tau_{i}}$, so $t\leq t_{i}+\Delta$. Furthermore, $\tau\notin P_{\tau_{j}}$, so $t+\Delta\geq t_{j}$. Furthermore, by Observation 6.2, $t_{j}+\Delta\geq t_{i}$.
 Thus, $t_{j}-\Delta\leq t\leq t_{i}+\Delta\leq t_{j}+2\Delta$. ∎
@@
 ∎
-8. Protocol Instantiation
-The algorithms thus far have assumed that every replica will eventually vote on every transaction.
-Any faulty replica, however, could exclude a transaction from its output, leaving the transaction (and the protocol) permanently stuck.
-The main idea to resolve this difficulty is that if a replica refuses to vote on a transaction, then that replica must be faulty.
-If a replica is faulty, then it suffices for the rest of the replicas to make up a vote on its behalf.
-We give here an example method for instantiating our sequencing protocol on top of existing protocol components, although this is far from the only method.
-As a base layer, we require some consensus protocol, such as (yin:hotstuff; lokhava:stellar; castro1999practical; miller2016honey), that proceeds in rounds.
-At any time, every replica can submit a continuation of its previous ordering; that is, an ordered list of transactions that extends its ordering vote.
-The consensus protocol then comes to consensus on a set of received ordering votes.
-The Streaming Ranked Vote algorithm is run (by every replica) on the set of received votes after every new set of votes is finalized by the consensus protocol.
-Assuming correctness of the consensus protocol, every replica will run the algorithm on the same input, and thereby produce the same output ordering.
-We assume that every replica has a registered public key.
-Every vote continuation must be cryptographically signed by the associated secret key.
-Furthermore, every continuation must be accompanied by the hash of the previously submitted continuation, so that if one continuation vote from a replica is lost (perhaps by malicious activity of consensus participants) the replica’s ordering cannot be manipulated.
-We make a network synchrony assumption, akin to Assumption 7.2.
-If a transaction appears first in some replica’s vote in round $T$ of the consensus protocol, then we assume that there is some known bound $k$ such that every other honest replica will have time to receive and submit a vote on the transaction before the end of round $T+k$.
-As such, after every round $r$, the protocol computes the set of all transactions that have not received votes from every replica but were voted on by some replica at or before round $r-k$.
-For each of these transactions $\tau$, the protocol appends a vote for $\tau$ at the end of a replica’s current vote sequence if $\tau$ is not already present.
-When appending multiple transactions to a replica’s vote, transactions can be sorted deterministically (i.e.
-by hash).
-This protocol guarantees that every transaction will receive a vote from all replicas within at most $k$ rounds after it is first submitted to the protocol.
-In this setting, $k$ must be configured in advance by the protocol.
-The parameter must be set so that every honest replica is able to submit a vote on a transaction within $k$ rounds.
-Naturally, the choice of $k$ depends on network conditions the censorship-resistance (or “chain-quality” (garay2015bitcoin)) of the consensus protocol.
-As an example alternative, one could also use the protocol in kelkar2020order, which is based on atomic broadcast and a “Set Byzantine Agreement” primitive.
-This protocol waits until all honest replicas are able to add a transaction $\tau$ to their ordering votes, and then invokes a protocol to agree on the set of replicas $U_{\tau}$ which have the transaction in their ordering votes.
-This set may not be the set of all replicas, but includes all honest replicas.
-To use our ordering protocol on this consensus protocol, it suffices to make up a vote for the (presumed faulty) replicas not in $U_{\tau}\cap U_{\tau^{\prime}}$.
-Replicas in $U_{\tau}\setminus U_{\tau^{\prime}}$ can be presumed to vote $\tau\prec\tau^{\prime}$ and vice versa, and replicas not in $U_{\tau}\cup U_{\tau^{\prime}}$ can be presumed to vote (arbitrarily) by, e.g., comparing deterministic hashes of $\tau$ and $\tau^{\prime}$.
+7. Related Work
+7.1. Order-Fairness
+Closest to this work are Aequitas (kelkar2020order) and Themis (themis), which propose the definition of $\gamma$-batch-order-fairness and instantiate ordering protocols based on aggregating ordering votes into a total order.
+Aequitas achieves $(\gamma,\frac{f}{n})$-minimal-batch-order-fairness for a fixed choice of $\gamma$, but is not asymptotically live.
+aequitaspermissionless instantiate Aequitas in a setting where the number of replicas is not known, and similarly points out the connection between this streaming problem and the classical ranking aggregation problem.
+§3.3 gives a more detailed comparison to these works.
+cachin2022quick give a notion of “$\kappa$-differential-order-fairness;” Theorem D.1 of (themis) shows equivalence of this definition and $\gamma$-batch-order-fairness, for a choice of $\gamma$ (as a function of $\kappa$).
+Our definition of $(\gamma,\delta)$-minimal-batch-order-fairness (Definition 3.4) strengthens the (vacuously satisfiable) $\gamma$-batch-order-fairness definition used in prior work (see §3 for details).
+Concurrently with this work, vafadar2023condorcet observe that an adversary can influence the output ordering of many protocols (including that of kelkar2020order) by sending extra transactions (which can force batches to be combined together).
+This eliminates the constraints on the output that Definition 3.1 gives (using an example akin to Example 10.2).
+Lemma 3.3 identifies a similar problem.
+This observation is a corollary of the fact that the protocol of kelkar2020order (as well as the Ranked Pairs algorithm studied here) does not satisfy the Irrelevance of Independent Alternatives axiom (arrow1950difficulty).
+vafadar2023condorcet also propose using Ranked Pairs to sort transactions within batches (as, e.g., those output by a protocol like Themis (themis)).
+However, no modification to the ordering of transactions within a batch output in Aequitas (kelkar2020order) can give the Ranked Pairs ordering, because Ranked Pairs might need to interleave batches that Aequitas considers incomparable (as demonstrated in §11).
+This strategy also cannot mitigate Aequitas’s lack of asymptotic liveness.
+This mitigation, if applied to Aequitas, is akin to dropping all edges of weight less than $\gamma$ before running Ranked Pairs.
+This strategy, therefore, would achieve $(\gamma^{\prime},\frac{f}{n})$-minimal-batch-order-fairness for all $\gamma^{\prime}\geq\gamma$ (a weaker guarantee than what our protocol provides).
+Also concurrently with this work, kiayias2024ordering propose a different method for ordering transactions within batches (but rely on a specific choice of a parameter $\gamma$ to determine those batches).
+A version of that protocol adds timestamps on reported orderings to improve liveness.
+As discussed in §3.3, the batches in Themis (themis) are defined slightly differently than in Definition 3.1.
+Themis, when computing a graph of ordering dependencies, adds a directed edge between every pair of vertices, choosing the direction of an edge by majority vote (ignoring the parameter $\gamma$, which is considered only on edges adjacent to transactions that some replicas have not yet reported observing).
+In so doing, it merges together any two batches unless, for any transaction $\tau$ in one and $\tau^{\prime}$ in the other, a majority of replicas vote $\tau\prec\tau^{\prime}$ (which bypasses the counterexample of §11).
+Additionally, Themis’s (themis) round-based sequencing protocol presents a different problem insurmountable by any process like the one suggested by vafadar2023condorcet that sorts batches separately.
+Namely, Themis may finalize the composition of a batch before has enough information to guarantee that, according to the Ranked Pairs ordering, all transactions in the batch must come before all those not in the batch (and not already in the output).
+Thus, it may output two batches in sequence, such that running Ranked Pairs on each batch separately produces a different output than running Ranked Pairs on the union of the two batches together.
+We give a example of this phenomenon in §12.
+zhang2020byzantine add timestamps to ordering votes and sorts transactions by median timestamp to provide a linearizability guarantee, which is incomparable to $(\gamma,\delta)$-minimal-batch-order-fairness.
+mamageishvili2023buying propose letting users pay to reduce their reported timestamps.
+7.2. Social Choice
+Classical work of social choice studies the problem of choosing an order between a fixed, finite set of candidates given a complete set of preferences from each voter (arrow1950difficulty).
+Prior work studies related models in a limited-information setting.
+Lu and Boutilier (lu2013multi), Ackerman et.
+al (ackerman2013elections), and Cullinan et.
+al (cullinan2014borda) study the setting where a social choice rule has incomplete access (a partial ordering) to a voter’s preferences over a finite set of candidates.
+Conitzer and Sandholm (conitzer2005communication) study the communication complexity of a variety of voting rules.
+Fain et.
+al (fain2019random) study a randomized ordering rule that requires only a constant number of queries to voter preferences.
+Fishburn (fishburn1970arrow) showed that Arrow’s impossibility theorem does not hold when the set of voters is infinite, although Kirman and Sondermann (kirman1972arrow) find “dictatorial sets” of voters.
+Grafe and Grafe (grafe1983arrow) extend these results to the case of infinitely many alternatives, subject to a continuity condition on the space of ranking preferences.
+Chichilnisky and Heal (chichilnisky1997social) and Efimov and Koshevoy (efimov1994topological) study the types of rules admissible in the infinite voters setting, given a topology on ordering preferences.
+7.3. Front-Running
+The study of order-fairness is often motivated by the goal of limiting an adversary’s ability to order transactions (“front-running”) in public blockchains (as in, e.g., daian2020flash).
+Additional approaches to this problem include commiting to an ordering before any replica can know transaction contents, using threshold encryption or commit-reveal schemes (malkhi2022maximal; clineclockwork; zhang2022flash).
+Li et.
+al (li2023transaction) study the problem of verifiably computing an ordering within a trusted hardware enclave.
+These could be applied on top of a streaming ordering algorithm.
+Kavousi et.
+al (kavousi2023blindperm) randomly shuffle blocks of encrypted transactions.
+Some decentralized exchanges (ramseyer2023speedex; cowswapproblem; penumbraswap) process transactions in unordered batches, eliminating the need for an ordering within a block (but not eliminating all order manipulations (zhangcomputation)).
+ferreira2022credible and li2023mev construct sequencing rules specific to decentralized exchanges.
 8. Conclusion